### 04 Apr 2026 v1

# April 2026

Abstract
We study the role of commutativity in the representation of relativistic locality for localization observables of relativistic quantum systems in Minkowski spacetime. A celebrated no-go theorem by Halvorson and Clifton shows that commutativity of localization effects associated with causally separated regions is incompatible with other apparently natural assumptions on spatial localization. Commutativity is in particular assumed to be the mathematical representation of locality in the Araki-Haag-Kastler formulation of QFT. This raises the question of whether commutativity follows from more elementary locality principles of Quantum Theory. Adopting Busch’s operational analysis in terms of no-signaling and relativistic consistency, we argue that, for a particle-like system, the commutativity requirement is not a consequence of these principles. Under a natural local detectability principle, elementary localization observables are not localized in arbitrarily small spacetime neighborhoods of the corresponding spatial regions, but rather in regions containing the whole rest space (a Cauchy surface) where the measurement is performed. This is a consequence of the very nature of particle which is assumed to localize in a unique place on a rest space completely filled with ideal detectors. To this regard there is no direct conflict with Araki-Haag-Kastler formulation of local quantum physics. On the other hand we also shows that commutativity and localization may coexists if referring to less idealized localization procedures. Indeed, we introduce conditional localization POVMs associated with bounded spatial regions interpreted as laboratories. These observables, due to a result in quantum information theory known as gentle measurement lemma, describe conditional localization probabilities. In principle, these observables satisfy commutativity when associated with causally separated laboratories. As a consequence they could be represented by local observables in the sense of Araki-Haag-Kastler. Explicit examples of these local observables satisfying a commutative requirement will be presented is a forthcoming paper within the approach of local QFT.
1 Introduction
1.1 Localization, Causality, Commutativity
Spatial localization in relativistic quantum theory has long been a delicate issue, lying at the intersection of quantum measurement theory, spacetime causality, and the very notion of particle.
In non-relativistic quantum mechanics, localization is naturally described by projection-valued measures and, correspondingly, by selfadjoint position operators. This is a case of sharp localization, since the projectors $P(\Delta)$ of the spectral measure are labelled by arbitrarily small regions $\Delta$ of the rest space $\Sigma$ of an observer, and we can prepare states $\rho$ for which the probability of detecting a particle a given such bounded region is $tr(\rho P(\Delta))=1$.
In the relativistic setting, however, this scheme encounters severe difficulties. The best-known example is the Newton-Wigner construction, which provides a mathematically neat notion of sharp position for relativistic particles, but whose physical interpretation has long been debated. In particular, Newton-Wigner localization was proposed as a possible tool for reinterpreting localization in (free) quantum field theory so as to avoid some of the puzzling consequences related to the Reeh-Schlieder property [Fle00]. Nevertheless, it was later shown that the counterintuitive features implied by the Reeh-Schlieder theorem persist also in this reformulation [Hal01].
From the one-particle viewpoint, a further serious issue is raised by Hegerfeldt’s celebrated analysis [Heg74, Heg85]. If energy, i.e., the selfadjoint generator of time translations is positive (or bounded from below), sharp localization described by projection-valued measures gives rise to causality problems. In modern terms, and especially in the formulation later clarified by Castrigiano, these problems concern the incompatibility of sharp localization with a natural requirement of causal (i.e. non-superluminal) propagation for localization probabilities. This strongly suggests that, in relativistic quantum theory, sharp notions of localization should be abandoned in favour of unsharp ones described by POVMs. In this case, effects, that is, positive operators bounded by the identity, $0\leq A(\Delta)\leq I$, rather than orthogonal projectors, are still associated with arbitrarily small regions $\Delta$ in the rest space $\Sigma$ of an observer, but the probability of finding the particle in a given bounded region is strictly $tr(\rho A(\Delta))<1$ for every state $\rho$, even though sharp states can be approximated arbitrarily well.
Recent works [Cas17, Mor23, Cas24, DRMo24, Cas25, CDM26] have shown that this strategy is in fact viable. A broad family of relativistic spatial localization observables has been constructed and analyzed for various elementary systems, with full compatibility with Poincaré covariance and with natural causal requirements, especially in the unsharp setting when the reqirement on positive energy is assumed. These developments indicate that Hegerfeldt’s issue does not rule out unsharp localization under the requirement of energy positivity, but rather rules out an excessively rigid, projection-based notion of localization.
There is, however, another and subtler problem, which concerns not the propagation of localization probabilities in time, but rather the representation of relativistic locality itself in measurement theory. A central result in this direction is the no-go theorem by Halvorson and Clifton [HaCl02], which may be regarded as the culmination of a line of investigation initiated by Schlieder [Sch71] and Jancewicz [Jan77], sharpened by Malament [Mal96] for projector-valued localization, and later extended by Busch [Bus99]. Roughly speaking, Halvorson and Clifton proved that, under a family of assumptions that appear very general and natural for a relativistic localization scheme, the localization effects (elements of an alleged localization POVM) must be trivial. In this sense, the theorem seems to suggest a deep tension between relativistic localization and relativistic locality. We stress that the theorem does not assume a particular definition of the considered quantum system which, in principle, could be formalized as a Wigner elementary particle, a composite system or even a quantum field. The assumptions regard additivity, translation covariance, positivity of energy, and a microcausality requirement expressed by commutativity of localization effects associated with suitably ”separated” regions $\Delta,\Delta^{\prime}$:
$A(\Delta),A(\Delta^{\prime})]=0\>.$
Actually it is not even supposed that a whole unitary representation of the Poincaré group exists. The relativistic perspective comes into the play just to physically justify commutativity of localization effects associated to ”separated” regions, when interpreted as causally separated regions in relativistic sense.
The crucial point is that the commutativity requirement entering the Halvorson-Clifton theorem is usually regarded as self-evident from the perspective of local quantum physics, especially within the Araki-Haag-Kastler framework [Haa96, Ara09], where locality is encoded precisely by commutativity of observables associated with causally separated spacetime regions. If the spatial region $\Delta$ of a restapace $\Sigma$ where a particle is found at a certain time $t$ belongs to the spacetime region $\cal O$} where the localization-observable $A(\Delta)$ is localized (in AHK sense), then the commutativity assumption would indeed appear unavoidable. From this standpoint Halvorson-Clifton’s result would represent an insurmountable obstruction to any notion of position of a quantum system at given time in a reference frame. However, this line of reasoning already presupposes a specific mathematical representation of locality. It is therefore natural to investigate the interplay of commutativity and causally separation of the localization region, independently of the assumption that localization observables $A(\Delta)$ belong to AHK local algebras ${{\cal A}}({\cal O})$, is strict sense, attached to arbitrarily small spacetime neighborhoods $\cal O$ of the corresponding spatial regions: ${\cal O}\supset\Delta$. A first step in this direction was taken by Busch [Bus99], building on a technical result with Singh [BuSi98], by analyzing the relation between commutativity and locality directly within the general theory of quantum measurement. In Busch’s approach an observable is localized in a region $\cal O$} of spacetime if it can be measured with instruments which work therein. This is also in the spirit of AHK structure, but contrarily to that framework, no further requirement about the structure (*, $C^{*}$, von Neumann algebra) is assumed for the family of observables localized in a region of spacetime. Similarly, no a priori requirements for families of observables associated to distinct regions of spacetime are supposed. In that setting, locality is formulated through operational principles such as the no-signaling condition and, in a stronger form, what is often called relativistic consistency [Bec20], namely invariance of the statistics of suitable consecutive measurements under interchange of their temporal order when the measurements take place in causally separated regions. For broad classes of measurement procedures – in particular for Lüders measurements and related efficient schemes – these principles imply commutativity of the relevant effects. This analysis is mathematically precise and conceptually important because it does not assume locality in the algebraic form of microcausality (i.e. commutativity), but rather attempts to derive commutativity from independent operational requirements.
The main purpose of this paper is to investigate whether this route can actually justify the commutativity requirement for relativistic spatial localization observables. The answer we shall propose is, in general, negative. To this end, following Busch’s approach, we adopt in particular a natural local detectability principle which asserts that if instruments work in an open spacetime region $\cal O$} and they are used to detect a particle in a spatial region $\Delta$, then $\Delta\subset{\cal O}$.
The above negative answer is not due to a failure of no-signaling or relativistic consistency, but rather in the physical meaning of localization for a quantum system which we define to be a particle. Let us assume that a localization observable is meant to describe a system with the propensity to be found in a unique place when an ideal position measurement is performed on a whole rest space. Then the elementary observable associated with detection in a region $\Delta$ is not naturally localized in an arbitrarily small spacetime neighborhood of $\Delta$. On the contrary, once one adopts the above stated local detectability principle, as we shall prove in Section 4.2, one is led to the conclusion that the elementary observable $\{A(\Delta),I-A(\Delta)\}$ is localized, in Busch sense, in a region that must contain the whole rest space where the global localization experiment is performed. As a consequence, two elementary localization observables associated with spatially disjoint regions are not localized in causally separated spacetime regions, and therefore the no-signaling and relativistic-consistency arguments leading to commutativity cannot even be triggered. In this sense, the commutativity requirement for localization effects does not follow from those more elementary locality principles, since the very hypotheses for employing these principles are not satisfied. This conclusion is physically tied to the particle character of the systems under consideration. A particle is assumed not to localize simultaneously in several (causally) disjoint regions in an ideal exhaustive detection experiment, but rather in a unique place. If one measures whether the particle is found in a region $\Delta$ within a complete position measurement on a rest space $\Sigma$, then one automatically obtains information about the complementary region $\Sigma\setminus\Delta$. From this viewpoint, the associated elementary observable is not confined to $\Delta$ alone: it carries information about the global detection arrangement on $\Sigma$. The failure of the commutativity argument is therefore not an exotic mathematical pathology, but rather a direct consequence of the physical meaning attributed to localization of a particle. If we do not accept this, we cannot accept the very notion of a particle.
These arguments provide a motivation, for asserting that localization effects, for particles, do not belong to local algebras in the sense of AHK formulation, in the case where they are parts of a POVM which fills the whole rest space of a reference frame (a Cauchy surface).
The second aim of the paper is to indicate a possible way to make compatible spatial localization and commutativity in more realistic localization experiments.
We introduce a notion of conditional localization associated with a bounded spatial region $\Delta_{0}$, interpreted as the spatial part of a laboratory. The idea is to construct, starting from a general localization observable defined on a full rest space, a new POVM normalized on $\Delta_{0}$ which describes the conditional probability of finding the system in a subregion $\Delta\subset\Delta_{0}$, given that the system has been detected in $\Delta_{0}$. The mathematical possibility of such construction relies upon a known result in quantum information theory called gentle measurement lemma. It permits to prove that a specific type of POVMs we introduce in a laboratory normalized on bonded spatial region $\Delta_{0}$, measure the conditional probability to find a particle in $\Delta\subset\Delta_{0}$ if we know that the system was detected in $\Delta_{0}$. This is all the more true the closer the probability of finding the particle in $\Delta_{0}$ is to $1$. These conditional POVMs are intrinsically local to the laboratory and, unlike the original localization observables on the whole rest space, are not affected by the same problem that prevents the use of no-signaling or relativistic-consistency arguments. We shall argue that pairs of such conditional POVMs associated with causally separated laboratories may, in principle, satisfy commutativity and belong to local algebras concentrated around the laboratory region in the AHK sense. This suggests that conditional localization may provide a mathematically and physically appropriate notion of local position measurement in relativistic quantum theory. Explicit examples of these local observables satisfying a commutative requirement and constructed in terms of local energies will be presented is a forthcoming paper within the approach of local QFT.
1.2 Structure of the work
The paper is organized as follows. In Section 2 we review the notion of relativistic spatial localization observable and the relevant causal conditions, and we recall the Halvorson-Clifton no-go theorem for unsharp localization systems. We also discuss the relation of the problem with the Araki-Haag-Kastler framework. In Section 3 we summarize Busch’s operational analysis of locality in terms of no-signaling and relativistic consistency, including the relevant results connecting these principles with commutativity of effects. In Section 4 we apply that analysis to relativistic localization observables and prove that, under a natural local detectability principle and for systems with particle-like localization behaviour, commutativity of causally separated localization effects is not implied by those operational locality principles. In Section 5 we introduce conditional localization POVMs in bounded laboratories, study their basic mathematical properties, and discuss why they may evade the previous problem. The paper closes with conclusions and an outlook toward a subsequent work, where explicit examples of localization observables and conditional local position observables will be constructed within a local quantum-field-theoretic framework.
1.3 Notions and notation
We assume $c=1$, $\hbar=1$ throughout the rest of this paper, the notation $A\subset B$ allows the case $A=B$. All Hilbert spaces are complex and symmetric operators are densely defined by definition.
Minkowski spacetime. A four-dimensional real affine space whose space of translations $\mathsf{V}$ is equipped with a non-degenerate symmetric bilinear form $g$ with signature (,+,+,+)(-,+,+,+) is called the Minkowski spacetime $\mathbb{M}$}. The points of $\mathbb{M}$} are called events, and $g$ is called the Minkowski metric. A vector $v\in\mathsf{V}$ is spacelike if $g(v,v)>0$ or $v=0$. It is causal if $g(v,v)\leq 0$ and $v\neq 0$. A causal vector $v$ is timelike if $g(v,v)<0$, or lightlike if $g(v,v)=0$. Smooth curves are classified analogously according to their tangent vectors.
A set $\Lambda\subset{\mathbb{M}}$ is achronal if pqp-q cannot be timelike for $p,q\in\Lambda$.
The set of timelike vectors is an open cone consisting of two disjoint open connected components. A choice of one of them $\mathsf{V}_{+}$ defines a time orientation of $\mathbb{M}$}. The latter is henceforth assumed to be time oriented: $\mathsf{V}_{+}\subset\mathsf{V}$ is the open cone of future-directed timelike vectors. $\overline{\mathsf{V}_{+}}\setminus\{0\}$ is the cone of future-directed causal vectors. Notice that if $v\in\overline{\mathsf{V}_{+}}$ and $u$ is causal, then $u\in\overline{\mathsf{V}_{+}}$ if and only if $g(u,v)\leq 0$. We finally define ${\mathsf{T}}_{+}:=\{u\in\mathsf{V}_{+}\>|\>g(u,u)=-1\}$.
If $A\subset{\mathbb{M}}$, the causal future $J^{+}(A):=\{p\in{\mathbb{M}}\>|\>p-q\in\overline{\mathsf{V}_{+}}\quad\mbox{for some $q\in A$}\}$ represents the events of $\mathbb{M}$} in the future of $A$ that can be physically influenced by $A$. The causal past $J^{-}(A):=\{p\in{\mathbb{M}}\>|\>q-p\in\overline{\mathsf{V}_{+}}\quad\mbox{for some $q\in A$}\}$ is defined symmetrically.
$A,B\subset{\mathbb{M}}$ are said to be causally separated if $(J^{+}(A)\cup J^{-}(A))\cap B=\varnothing$ (which is equivalent to $(J^{+}(B)\cup J^{-}(B))\cap A=\varnothing$).
If $R\subset{\mathbb{M}}$, its causal complement and its causal completion, respectively, are
$R^{\perp\!\!\!\!\perp}:={\mathbb{M}}\setminus((J^{+}(R)\cup J^{-}(R)))\qquad\mbox{and}\qquad(R^{{\perp\!\!\!\!\perp}})^{{\perp\!\!\!\!\perp}}\>.$ (1)
It is easy to prove that $R\subset(R^{{\perp\!\!\!\!\perp}})^{{\perp\!\!\!\!\perp}}$. If $R\subset{\mathbb{M}}$, $(R^{{\perp\!\!\!\!\perp}})^{{\perp\!\!\!\!\perp}}$ turns out to consist of the points $p\in{\mathbb{M}}$ such that every causal straight line passing through $p$ meets $R$ somewhere [^1] This is equivalent, in $\mathbb{M}$}, to the set of points $p$ such that every inextensible causal curve passing through $p$ also meets $R$ somewhere. This latter set, in a generic spacetime $M$, is also called the domain of dependence of $R$ when $R$ is achronal..
A Minkowskian reference frame – physically representing an inertial reference frame or an observer – is a unit timelike vector $n\in{\mathsf{T}}_{+}$.
A rest space of a Minkowski reference frame $u\in{\mathsf{T}}_{+}$ is an affine $3$-plane $\Sigma\subset{\mathbb{M}}$ $g$-normal to $u$, written $u\perp\Sigma$. If $o\in{\mathbb{M}}$ is a given origin, the family of rest spaces of $u\in{\mathsf{T}}_{+}$ is labeled by the time at which they occur in the said reference frame. Every spacelike affine $3$-plane is the rest space of a Minkowskian reference frame, and is said to be adapted to $\Sigma$, at some time.
$d\Sigma$ denotes the natural translationally invariant Borel measure on the rest space $\Sigma$ of $u$: the Lebesgue measure on $\Sigma$ induced by the metric. $\mathscr{B}(\Sigma)$ denotes the family of Borel subsets of $\Sigma$. We use the notation $|\Delta|=\int_{\Sigma}\chi_{\Delta}(x)d\Sigma(x)$ if $\Delta\in\mathscr{B}(\Sigma)$.
The Lie group of metric-preserving affine maps $h:{\mathbb{M}}\to{\mathbb{M}}$ is known as the Poincaré group $IO(1,3)$. The Lie subgroup of affine maps that also preserve the time orientation is called the orthochronous Poincarè group $IO(1,3)_{+}$. The subgroup of $IO(1,3)_{+}$ that leaves fixed an arbitrarily chosen origin [^2] Different choices of $o$ give rise to isomorphic definitions and the component $\Lambda$ of an element of $IO(1,3)_{+}$ does not depend on the choice of $o$. $o\in{\mathbb{M}}$ is the orthochronous Lorentz group $O(1,3)_{+}$. Elements $h\in IO(1,3)_{+}$ are in one-to-one correspondence with pairs $(\Lambda,v)$ where $v\in\mathsf{V}$ and $\Lambda\in O(1,3)_{+}$, and the action of $h$ on a point $q\in{\mathbb{M}}$ is
$(\Lambda,v)q=o+v+\Lambda(q-o)\>.$ (2)
The subgroup of $IO(1,3)_{+}$ known as the proper orthochronous Poincaré group $ISO(1,3)_{+}$ is obtained by replacing $O(1,3)_{+}$ with the proper orthochronous Lorentz group $SO(1,3)_{+}$. The latter, representing $O(1,3)_{+}$ in a Minkowskian coordinate system as above, is obtained by restricting to Lorentz matrices with $\det\Lambda>0$.
For some physical systems, such as fermions, Poincaré invariance is implemented in terms of projective unitary representations $V:ISO(1,3)_{+}\to{\mathfrak{B}}({{\cal H}})$ such that they become genuinely unitary when viewed as representations of the universal covering of the proper orthochronous Poincarè group $\widetilde{ISO(1,3)_{+}}$. In other words (see e.g. [Mor19]), the representation $U:=V\circ\pi:\widetilde{ISO(1,3)_{+}}\to{\mathfrak{B}}({{\cal H}})$ is unitary if $\pi:\widetilde{ISO(1,3)_{+}}\to ISO(1,3)_{+}$ is the canonical projection homomorphism. This result is obviously valid even if the initial representation $V$ is genuinely unitary. As a consequence, we sometimes replace $ISO(1,3)_{+}$ with $\widetilde{ISO(1,3)_{+}}$ in some definitions in order to cover a larger class of physical situations.
Quantum observables in Hilbert space. ${\mathfrak{B}}({{\cal H}})$ denotes the $C^{*}$-algebra of bounded operators $A:{{\cal H}}\to{{\cal H}}$. If $\in{\mathfrak{B}}({{\cal H}}), ABA\geq B$ (equivalently $B\leq A$) means that $A-B\geq 0$; in turn, $A\geq 0$ means that $\langle\psi|A\psi\rangle\geq 0$ if $\psi\in{{\cal H}}$. In that case we say that the operator $A$ is positive.
[…]
Generally, mixed states $\rho$ are trace-class operators $\rho\in{\mathfrak{B}}_{1}({{\cal H}})$, positive ($\rho\geq 0$), and normalized ($\mathrm{tr}\rho=1$). The convex body of states will be denoted by ${\mathsf{S}}({{\cal H}})$. A special case of states is given by one-dimensional projectors $\rho=|\psi\rangle\langle\psi|$ for unit vectors [^3] These are pure states, i.e., extremal elements in the space of states if the von Neumann algebra of observables is the whole ${\mathfrak{B}}({{\cal H}})$., $\psi\in{{\cal H}}$.
For a state $\rho\in{\mathsf{S}}({{\cal H}})$, $tr(E(B)\rho)=tr(\rho E(B))$ is interpreted as the probability of obtaining an outcome in $B$ when the system is in the state $\rho$.
If $\Sigma({\mathbb{R}})$ is the Borel $\sigma$-algebra $\mathscr{B}({\mathbb{R}})$ and all the effects $E(B)$ are orthogonal projectors, we have a standard Projector-Valued Measure (PVM). As is well known, every PVM is in one-to-one correspondence with a selfadjoint operator $\hat{E}=\int_{{\mathbb{R}}}xE(dx)$ through the spectral theorem. In this sense, a POVM is a generalized observable.
In the special case where $\Sigma(X)$ is the power set of $X:=\{1,\ldots,N\}$, a POVM on $X$ is completely determined by the special effects $E_{j}:=E(\{j\})$ with $j\in\{1,\ldots,N\}$. As a general reference textbook on this mathematical technology applied to physics, we suggest [BLPY16]. References on general spectral theory applied to physics that we shall use are [Mor18, Mor19].
2 Halvorson–Clifton’s no-go result for unsharp localizations
In what follows, we first introduce some types of spatial localization, based on the POVM formalism, which are compatible with Hegerfeldt’s locality constraint. We shall simultaneously review Hegerfeldt’s causality conditions in the modern presentation due to Castrigiano. Afterwards, we shall present the Halvorson–Clifton no-go result [HaCl02]. Finally, we shall focus on the crucial commutativity hypothesis assumed by Halvorson and Clifton from the standpoint of Busch’s analysis of locality [Bus99] and of more recent developments on the subject, especially those collected in Beck’s paper [Bec20].
2.1 Types of unsharp spatial localization
[…]
Definition 2.1 (Unsharp Localization System):
An unsharp localization system in Minkowski spacetime $\mathbb{M}$} is a quadruple $({{\cal H}},{\cal R}_{0},A_{0},U_{0})$ where:
  • (a)
[…]
  • (b)
    $A_{0}:\mathcal{R}_{0}\ni\Delta\mapsto A_{0}(\Delta)\in{\mathfrak{B}}({{\cal H}})$ is an effect-valued map ($0\leq A_{0}(\Delta)\leq I$), with $\mathcal{R}_{0}$ a family of bounded sets of a foliation of parallel spacelike 3-planes in $\mathbb{M}$};
  • (c)
    $U_{0}:\mathsf{V}\ni a\mapsto U_{0a}\in{\mathfrak{B}}({{\cal H}})$ is a strongly continuous unitary representation of the translation space $\mathsf{V}$ of the affine space $\mathbb{M}$}. $\blacksquare$
The above structure is sufficient for stating the most basic results. However, for the development of our discussion we shall need a more elaborate construction, which specializes Definition 2.1 and has also been used in recent literature. From a physical perspective, the mathematical machinery introduced in Definition 2.1 does not seem sufficient to capture the physical requirements embodied in an (idealized) exhaustive detection experiment. If we want to interrogate a quantum system about its position in space, we should force it to take a position in space. Assuming that a detection is an interaction with an instrument, this means that we have to fill the entire space with detectors. Within this setting, the system localizes somewhere at the given time. If detectors are mathematically represented by effects $A_{0}(\Delta)$, we are forced to assume that either infinitely extended regions $\Delta$ are permitted, or the map $\Delta\mapsto A_{0}(\Delta)$ is $\sigma$-additive (and not simply additive), or both, in order to make contact with the standard formulation of probability theory. Another observation is that the relativity principle suggests covariance with respect to the whole orthochronous Poincaré group. According to these ideas, we can state the following general definition of a relativistic spatial localization observable, readapted [^4] In those references the POVMs on the various $\Sigma$ were defined independently with a further coherence condition, and an absolute continuity condition was added in [DRMo24]. from [Mor23, DRMo24]. This notion, in a slightly simplified version, was already introduced and deeply analyzed for the first time by Castrigiano (see [Cas17] and references therein) under the name of Poincaré covariant POL.
Definition 2.2 (Relativistic Spatial Localization Observable):
A relativistic spatial localization observable is a quadruple $({{\cal H}},{\cal R},A,U)$ where
  • (a)
    ${\cal R}=\cup\{\mathscr{B}(\Sigma)\>|\>\Sigma\subset{\mathbb{M}}\>\mbox{spacelike $3$-plane}\}$ where $\mathscr{B}(\Sigma)$ is the Borel $\sigma$-algebra on $\Sigma$;
  • (b)
[…]
  • (c)
    $U:IO(1,3)_{+}\to{\mathfrak{B}}({{\cal H}})$ is a strongly-continuous unitary representation of the the orthochronous Poincaré group;
  • (d)
    AA is $U$-covariant, i.e., $U_{g}A^{v}(\Delta)U_{g}^{-1}=A^{gv}(g\Delta)$ for every $g\in{IO(1,3)_{+}}$ and $\Delta\in{\cal R}$. where ${\mathsf{S}}\ni v\mapsto gv\in{\mathsf{S}}$ denotes the action of $IO(1,3)_{+}$} on the tensors in $\mathsf{S}$}.
In (a), possibly ${\mathsf{S}}=\{1\}$, so that $A$ includes a unique element denoted by the same symbol $A$. In this case the relativistic spatial localization observable is said to be scalar.
In (c), $U$ may be generalized to a strongly-continuous unitary representation of the universal covering of the proper orthochronous Poincaré group $\widetilde{ISO(1,3)_{+}}$ provided $g\Delta$ and $gv$ in (d) are respectively replaced by $\pi(g)\Delta$ and $\pi(g)v$ – where $\pi:\widetilde{ISO(1,3)_{+}}\to{ISO(1,3)_{+}}$ is the canonical homomorphism projection.
[…]
At this stage, we can state, in modern language, the causality conditions implicit in Hegerfeldt’s original analysis [Heg74, Heg85] concerning the impossibility of superluminal propagation of detection probability.
Definition 2.3 (Castrigiano’s causality conditions):
A relativistic spatial localization observable $({{\cal H}},{\cal R},A,U)$ is causal if it satisfies the causal condition: for every $\Delta\in{\cal R}$, every spacelike $3$-plane $\Sigma$ and every $v\in{\mathsf{S}}$
  • ($C$)
       $A^{v}(\Delta)\leq A^{v}(\Delta_{\Sigma})\quad\mbox{where $\Delta_{\Sigma}:=\Sigma\cap(J^{+}(\Delta)\cup J^{-}(\Delta))$ provided $\Delta_{\Sigma}\in{\cal R}$.}$
A relativistic spatial localization observable $({{\cal H}},{\cal R},A,U)$ satisfies the causal time condition with respect to a Minkowskian reference frame $u\in{\mathsf{T}}_{+}$ [^5] The original notion of CT [Cas17] does not assume Poincaré covariance and refers to a preferred notion of time evolution. Assuming covariance, our definition immediately arises from the original one., if for every pair of spacelike $3$-planes $\Sigma_{0},\Sigma$ normal to $u$ and $\Delta\in{\cal R}$, with $\Delta\subset\Sigma_{0}$, and every $v\in{\mathsf{S}}$ it holds
  • (CT)
       $A^{v}(\Delta)\leq A^{v}(\Delta_{\Sigma})\quad\mbox{where $\Delta_{\Sigma}:=\Sigma\cap(J^{+}(\Delta)\cup J^{-}(\Delta))$ provided $\Delta_{\Sigma}\in{\cal R}$.}$ $\blacksquare$
Evidently, CC $\Rightarrow$ CT, but the converse implication is generally false (see [Mor23]).
Remark 2.4:
  • (1)
    It turns out (9 Theorem [Cas17]) that every relativistic spatial localization observable defined on a separable Hilbert space666as is the case for the Hilbert spaces we meet in this work., merely in view of covariance with respect to the translation group of a spatial $3$-plane $\Sigma$, satisfies $A^{v}(\Delta)=0$, for $\Delta\in\mathscr{B}(\Sigma)$, if and only if $|\Delta|=0$. In particular, $\mathscr{B}(\Sigma)\ni\Delta\mapsto\langle\psi|A^{v}(\Delta)\psi\rangle$ is absolutely continuous with respect to the Lebesgue measure $d\Sigma$ on the $3$-plane $\Sigma$ for every given $\psi\in{{\cal H}}$.
  • (2)
    As a consequence of (1), every $A^{v}|_{\mathscr{B}(\Sigma)}$ can be uniquely completed to a POVM defined on the Lebesgue $\sigma$-algebra $\mathscr{L}(\Sigma)$.
  • (3)
[…]
We stress that CC includes the original causality condition discussed by Hegerfeldt in his celebrated works [Heg74, Heg85], which, in the modern Castrigiano setting, corresponds to CT. The latter was proved to be violated by Hegerfeldt for relativistic spatial localization observables described in terms of PVMs, the Newton-Wigner localization in particular. The puzzling result was established under the hypothis that the generators of timelike translations have spectrum bounded from below [Cas17, Mor23]. Violation of CT implies violation of CC. In summary, if assuming CC, we are forced to abandon every sharp notion of localization described by PVMs in order to comply with causality. The explicit examples of relativistic spatial localization observables presented in [Cas17, Mor23, DRMo24, Cas24, Cas25, CDM26] for various types of particles [^7] These relativistic spatial localization observables when associated to particles have positive selfadjoint generators of time translation. show that in fact Hegerfeldt’s causal issue can be considered harmless. That is because CC is satisfied provided one uses suitable unsharp localizations in terms of POVMs as in Definition 2.2.
The notion of unsharp localization as in Definition 2.2 has been extended even further to include notions of rest spaces $S$ which are less rigid than the planes $\Sigma$. They can be chosen as smooth Cauchy surfaces (especially spacelike) [DRMo24] or even (Lipschitz) maximal achronal sets of $\mathbb{M}$} [Cas25], not necessarily Cauchy surfaces. Adopting these more general settings, the (generalized) causal condition CC turns out to be equivalent [DRMo24, Cas25] to the normalization of the POVMs defined on all (generalized) rest spaces. As an important byproduct, it was recently proved in [CDM26] that the extension of the formalism to achronal sets leads to a representation of the so-called causal logic of spacetime in terms of effects, when dealing with massive bosons.
2.2 The no-go result by Halvorson and Clifton
Although Hegerfeldt’s causality issue seems to be solved by using a suitable unsharp notion of localization, a second type of causality issue remains. It concerns the representation of locality in terms of commuting operators associated with causally separated observables. The most modern version of this issue was formulated in the famous Halvorson-Clifton paper [HaCl02], extending Schlieder’s [Sch71], Jancewicz’ [Jan77], and Malament’s [Mal96] arguments (valid for sharp localization) to the case of unsharp localization.
Theorem 2.5 (Halvorson-Clifton):
Let $({{\cal H}},{\cal R}_{0},A_{0},U_{0})$ be an unsharp localization system in Minkowski spacetime satisfying the following properties:
  • 1.
    Additivity: ${\cal R}_{0}$ is closed with respect to finite unions of regions in the same 3-plane, and $A_{0}$ is additive thereon:
    $A_{0}(\Delta)+A_{0}(\Delta^{\prime})=A_{0}(\Delta\cup\Delta^{\prime})\quad\mbox{for $\Delta,\Delta^{\prime}\subset\Sigma$ with $\Delta\cap\Delta^{\prime}=\varnothing$ and $\Delta,\Delta^{\prime}\in{\cal R}_{0}$.}$
  • 2.
    Translation covariance: ${\cal R}_{0}$ is closed under translations of $\mathsf{V}$, namely $\Delta_{v}:=\Delta+{v}\in{\cal R}$ if ${v}\in\mathsf{V}$ and $\Delta\in{\cal R}_{0}$, and
    $U_{0a}A_{0}(\Delta)U^{-1}_{0v}=A_{0}(\Delta+{v}).$
  • 3.
    Energy bounded below: If ${u}\in\mathsf{V}_{+}$, the self-adjoint generator $H_{0u}$ of the strongly continuous unitary one-parameter group
    ${\mathbb{R}}\ni t\mapsto U_{0t{u}}=e^{-itH_{u}}$
    has spectrum bounded from below.
  • 4.
    Microcausality: Let $\Delta,\Delta^{\prime}\in{\cal R}_{0}$ lie in a common spacelike 3-plane $\Sigma\subset{\mathbb{M}}$, and suppose they are disjoint with strictly positive distance. If ${u}\in\mathsf{V}_{+}$, there exists $\delta>0$ such that
    $A_{0}(\Delta),A_{0}(\Delta^{\prime}+t{u})]=0\quad\mbox{for }0\leq t<\delta.$
Under hypotheses (1)-(4), it holds that $A_{0}(\Delta)=0$ for every $\Delta\in{\cal R}_{0}$.
Proof.
This is Theorem 2 in [HaCl02], whose proof appears in Appendix B therein. ∎
The result also applies to more general non-relativistic affine spacetimes, in which case a fifth hypothesis is necessary. We omit this hypothesis here since it is automatically satisfied in Minkowski spacetime. Nevertheless, there exist relativistic cases where the fifth hypothesis cannot be imposed [HaCl02].
Corollary 2.6:
No relativistic spatial localization observable according to Definition 2.2 satisfies condition (4) for elements $A^{v}(\Delta)$, with given $v\in\mathsf{V}$, if the selfadjoint generator of time translations through $U$ is bounded below.
Proof.
Relativistic spatial localization observables, when (fixing $v\in{\mathsf{S}}$ if any and) restricting to the bounded elements of ${\cal R}_{0}$ of any fixed foliation of parallel spacelike $3$-planes $\Sigma$, satisfy conditions (1)-(2) according to Definition 2.2. Hence the theorem applies if (3) is valid. As a consequence, a relativistic spatial localization observable cannot satisfy (4) in view of the normalization condition of the POVM defined by $A^{v}$ on $\Sigma$, which would be violated by the thesis of Theorem 2.5. ∎
The above corollary applies in particular to the relativistic spatial localization observables of Definition 2.2 constructed for relativistic particles [Mor23, Cas24, DRMo24, CDM26], and also to the causal POL of [Cas17] not constructed in terms of PVMs: for all these localization observables, condition (4) must be violated.
Halvorson-Clifton’s theorem seems to impose a severe limitation on any notion of spatial localization for quantum relativistic systems. This is because the first three hypotheses, which seem genuinely necessary for defining a notion of spatial localization, are in contradiction with the fourth one, which, on the other hand, seems to express an elementary locality requirement in the quantum framework. If one proved that hypothesis (4) cannot be relaxed because relaxing it would irremediably affect fundamental locality requirements, the conclusion would be that no notion of spatial localization makes sense.
A problematic point in this investigation is that commutativity of observables localized in causally separated regions is simply assumed as the basic description of locality in local quantum physics, in particular in the formalization due to Araki and Haag-Kastler (AHK) [Haa96, Ara09] (see the next section). A deeper scrutiny, independent of the AHK approach, seems to be necessary at this stage. In particular, where no requirement about the structure (*, $C^{*}$, von Neumann algebra) is assumed for the family of observables localized in a region of spacetime and no a priori requirements for families of observables associated to distinct regions of spacetime are supposed. This type of analysis was initiated by Busch [Bus99] and developed by further authors, relying on the general theory of quantum measurement and on general local causality requirements concerning consecutive measurement processes. This analysis will be summarized in Section 3.
2.3 Relation with the Araki-Haag-Kastler framework
[…]
$A_{1},A_{2}]=0\quad\mbox{if $A_{1}\in{\mathfrak{A}}({\cal O}_{1})$, $A_{2}\in{\mathfrak{A}}({\cal O}_{2})$ and ${\cal O}_{1}\cap(J^{+}({\cal O}_{2})\cup J^{+}({\cal O}_{2}))=\varnothing$.}$
[…]
There are a number of further requirements and fundamental results arising from the AHK approach. We shall, however, omit them all and stick to the elementary postulates above, since they are enough for our discussion. The reader may consult [Haa96, Ara09] for classic general treatments of the subject.
Most of the features of this approach can be generalized to the case where $\mathfrak{A}$} and every ${\mathfrak{A}}(\cal O)$ are unital *-algebras, in particular of operators in a given Hilbert space with a common invariant domain.
As already suggested, it seems natural to postulate that the self-adjoint operators $A^{v}(\Delta)$ and $A^{v^{\prime}}(\Delta^{\prime})$ belong to the local algebras associated with open spacetime subsets arbitrarily narrowed around $\Delta$ and $\Delta^{\prime}$, respectively. If these open sets are causally separated, adopting the AHK formulation, then $A^{v}(\Delta),A^{v^{\prime}}(\Delta^{\prime})]=0$. If $\Delta$ and $\Delta^{\prime}$ lie in a common $3$-space $\Sigma$ and the distance between them is strictly positive, it is easy to prove the precise version of commutativity stated in requirement (4) of Theorem 2.5 by means of this argument. We shall see in the next sections that the hypothesis that $A^{v}(\Delta)$ [^8] Actually the POVM $\{A^{v}(\Delta),I-A^{v}(\Delta)\}$. is localized in a spacetime region arbitrarily concentrated around $\Delta$ is, however, disputable when dealing with relativistic spatial localization observables of particles: if these structures exist and can be constructed within the AHK framework, the region of localization of the operators $A^{v}(\Delta)$ cannot be arbitrarily narrowed around $\Delta$.
As is well known (see [Bec20] in particular), this negative result also arises from a fundamental result in local QFT known as Reeh-Schlieder theorem [Ara09, Haa96], if one interprets $A^{v}(\Delta)$ as the mathematical representation of a detector localized in $\Delta$. If assuming that this operator is positive and that the vacuum state has zero expectation value (no particles stay in that state!) then one faces a mathematical contradiction. For this reason, in the rigorous local formulation of scattering theory in the Haag-Ruelle approach detectors are not assumed to be local observables but just quasi-local observables [Ara09, Haa96].
The AHK formulation makes it possible, in particular, to exploit modular theory [Ara09, Haa96] (in the one-particle space of a scalar free quantum field) when the net of observables is formulated in terms of von Neumann algebras. More recently, this perspective has inspired an alternative approach to localization, initiated by Lechner and de Oliveira [LedO26], based on lattice theory rather than on the space of effects, as in [CDM26]. The resulting formulation of localization gives rise to a fuzzy probability measure, which is not a genuine measure because it lacks additivity. Nevertheless, for separation scales larger than the Compton wavelength, the corresponding modular localization scheme becomes essentially additive and, remarkably, provides an approximation to the Newton–Wigner one.
We shall say some more on the AHK formulation in a forthcoming paper. Here we only stress that our analysis does not rely upon this formulation since we do not directly assume the AHK approach in this paper.
3 No-signaling, relativistic consistency, and commutativity
An important analysis of the validity of the commutativity hypothesis (4) in the Halvorson–Clifton theorem – relying solely on general principles of measurement theory, and thus independently of the AHK framework [Haa96, Ara09] – was presented by Busch in [Bus99] (drawing on a previous technical paper [BuSi98] co-authored with Singh), prior to the Halvorson–Clifton paper. In [HaCl02], Halvorson and Clifton justified their hypothesis (4) on the basis of Busch’s analysis.
The next three sections present and examine Busch’s argument and some subsequent developments on the subject, in particular those due to S. Gudder and collaborators [GuNa01, AGG02]. An updated and thorough discussion, including further relevant results and important remarks, appears in Section 2.2 of Beck’s comprehensive analysis developed in his PhD thesis [Bec20]. In Section 2.6 of Beck’s analysis, it is explicitly stated that local commutativity does not seem to be an entirely well-motivated requirement. We shall see that our results are consistent with this view.
3.1 Localization of (generalized) observables according to Busch
Busch-Singh’s paper [BuSi98] and, more precisely, Busch’s work [Bus99] focus on a version of Einstein causality called weak Einstein causality, also known as the no-signaling condition (NSC). A related causality condition mentioned by Busch is what we shall call the relativistic consistency condition (RCC), following Beck [Bec20]. To state them formally, we have to define the physical notion of localization for a generalized observable, starting from [Bus99]: “[…] we need to assume a physically meaningful association of observables with (bounded open) spacetime regions, in the sense that such observables can be measured by means of operations carried out within these regions”.
We stress that no further requirement about the structure (*, $C^{*}$, von Neumann, algebra) of the family of observables localized in a region is assumed contrarily to AHK approach. Similarly, no a priori requirements for families of observables associated to distinct regions of spacetime are supposed.
In this context, observables are POVMs, with PVMs as a special case. A selfadjoint operator is not directly an observable; the associated observable is its PVM. An effect $E$ of a POVM, in spite of being selfadjoint, is not necessarily interpreted as an observable in its own right. Indeed, the associated observable would be the PVM of $E$, which, however, plays no physical role in the general case. Nevertheless, $E$ defines an associated elementary observable, the POVM $\{E,I-E\}$. In the case of an orthogonal projector $P$, viewed as a test on the quantum system, its PVM $\{P,I-P\}$ coincides with the associated elementary observable.
Coming back to the notion of localized observable, we propose a slightly more general definition than Busch’s one (by relaxing the boundedness requirement), which is also technically more precise. It allows us to state NSC and RCC and to use them for both relativistic spatial localization observables and unsharp localization systems.
Definition 3.1:
A quantum observable $A$ (a POVM or a PVM) is said to be localized in an open (not necessarily bounded) spacetime region $\cal O$} if it can be measured by means of operations carried out within that region. $\blacksquare$
This definition has to be understood as a physical definition. From the mathematical viewpoint, we are simply defining a preferred relation in the Cartesian product of the set of spacetime regions and the set of observables.
When specializing to localization observables, we also explicitly adopt an apparently obvious physical principle [^9] Though the principle appears to be self-evident, the existence of entanglement phenomena such as those used in the original EPR argument could require a discussion. which we call the local detectability principle (LDP). The principle could already be stated for unsharp localization systems but, again, we want to include the case where the localization spatial set $\Delta$ is unbounded, as it may happen for relativistic spatial localization observables, for future convenience.
  • (LDP)
    Let $A(\Delta)$ be an effect either of an unsharp localization system or of a spatial localization observable. If the elementary POVM $\{A(\Delta),I-A(\Delta)\}$ is localized in an open region $\cal O$}, then ${\cal O}\supset\Delta$. $\blacksquare$
Let us illustrate the physical content of LDP. First of all detections are interactions which instruments: detectors. The used elementary POVM which appear above $\{A(\Delta),I-A(\Delta)\}$ physically represents an observable associated with a detector with two outcomes for a given state $\rho$.
  • 1: the detector clicks, with probability $tr(A(\Delta)\rho)$,
  • 0: the detector does not click, with probability $tr((I-A(\Delta))\rho)$.
Since phsical operations are not instantaneous and instruments occupy space larger than the region where the particle is detected, the detector operates in a spacetime region which, for instance, can be thought of as having the form ${\cal O}\equiv(-\epsilon,\epsilon)\times{\cal D}_{\Delta}$, for some $\epsilon>0$ where ${\cal D}_{\Delta}\subset\Sigma$ and the temporal coordinate refers to Minkowskian coordinates $(t,p)$ comoving with $n\perp\Sigma$. More generally, the temporal coordinate $t$ may vary in different intevals $(-\epsilon_{p},\epsilon^{\prime}_{p})$ where $p\in{\cal D}_{\Delta}$ and the maps $\Delta\ni p\mapsto\epsilon_{p},\epsilon^{\prime}_{p}\in(0,+\infty)$ are at least continuous. We can also consider different forms of the region $\cal O$} if also including (worldlines of) cables or the instruments to transmit information about where the particle is found. But it seems physically impossible not to assume that any region $\cal O$} where the elementary POVM $\{A(\Delta),I-A(\Delta)\}$ is localized does not include a region like the one represented above and thus $\Delta$ therein. This is what LDP states. In the case where $\Delta$ is unbounded, we can assume the use of an infinite family of detectors placed on corresponding bounded sets of a partition of $\Delta$. Saying that the system is detected in $\Delta$ means that one detector of the family clicks.
3.2 No-signaling condition and relativistic consistency condition
We now move on to discuss an elementary version of the no-signaling principle and a related principle that we shall call relativistic consistency. These physical principles are ways of making the locality principle precise in relativistic quantum theories [Bus99] without assuming a priori that locality corresponds to commutativity, as is supposed in the AHK framework. We need some technical definitions in order to state these principles. We shall deal with finite ($N<+\infty$) discrete POVMs because they are fully sufficient for the goals of this paper. However, all results could be extended to the case of infinite ($N=+\infty$) discrete POVMs.
Definition 3.2:
[…]
Under quite general hypotheses of quantum measurement theory [BLPY16], the theory of the post-measurement state is encapsulated in the following definition.
Definition 3.3:
The (selective) post-measurement state $\rho_{j}^{T}\in{\mathsf{S}}({{\cal H}})$ of an initial state $\rho\in{\mathsf{S}}({{\cal H}})$ corresponding to outcome $j$ of a finite discrete POVM $T$ as in Definition 3.2, and assuming that the probability $tr(\rho T_{j})$ of the outcome $j$ does not vanish, is
$\rho^{T}_{j}=\sum_{k\in I_{j}}\frac{K^{T}_{jk}\rho K^{T\dagger}_{jk}}{tr(\rho T_{j})},\qquad\text{with}\qquad T_{j}=\sum_{k\in I_{j}}{K^{T\dagger}_{jk}K^{T}_{jk}}\>.$ (3)
The operators $K^{T}_{jk}\in{\mathfrak{B}}({{\cal H}})$ are the Kraus operators associated with effect $T_{j}$. The set $I_{j}$ is here assumed to be finite [^10] It is possible to extend this formalism by relaxing these hypotheses [BLPY16] using suitable weak topologies; however, our setting is sufficient for the goals of this section..
The post-measurement state of a
non-selective measurement of $T$, given pre-measurement state $\rho\in{\mathfrak{B}}_{1}({{\cal H}})$, is (where $\mathrm{tr}(\rho T_{j})\rho^{T}_{j}:=0$ if $\mathrm{tr}(\rho T_{j})=0$)
$\rho^{T}:=\sum_{j=1}^{N}\mathrm{tr}(\rho T_{j})\rho^{T}_{j}=\sum_{j=1}^{N}\>\sum_{k\in I_{j}}{K^{T}_{jk}\rho K^{T\dagger}_{jk}}$ (4)
The map associating $\rho$ with $\rho^{T}_{j}$ or $\rho^{T}$ is a measurement (procedure) of $T$. It is said to be efficient if each $I_{j}$ consists of a unique element. $\blacksquare$
According to (4), the measurement of $T$ is non-selective if all measured systems of an initially prepared ensemble of identical systems in the state $\rho\in{\mathsf{S}}({{\cal H}})$ are finally collected together, regardless of the outcomes, yielding a new post-measurement state $\rho^{T}$.
We can now state our locality principles.
The no-signaling condition (NSC), taking Definition 3.1 into account, is stated as follows.
  • (NSC)
[…]
    The identity must hold
    $\mathrm{tr}(\rho S_{k})=\mathrm{tr}(\rho^{T}S_{k})\quad\mbox{for every state $\rho$ and every $k=1,\ldots,M$}$
    if ${\cal O}_{T}$ and ${\cal O}_{S}$ are causally separated. $\blacksquare$
In other words, the outcome statistics of $S$ must be identical whether we use $\rho^{T}$ or the original state $\rho$. Violation of NSC would allow one to infer, by observing only outcomes of $S$ in ${\cal O}_{S}$, whether a measurement of $T$ occurred in the causally separated region ${\cal O}_{T}$.
A related physical condition, which refers to selective measurements, is called the relativistic consistency condition (RCC):
  • (RCC)
[…]
    The identity must hold
    $\mathrm{tr}(\rho^{T}_{j}S_{i})\mathrm{tr}(\rho^{T}_{j})=\mathrm{tr}(\rho^{S}_{i}T_{j})\mathrm{tr}(\rho^{S}_{i})\quad\mbox{for every $\rho\in{\mathsf{S}}({{\cal H}})$ and $i=1,\ldots,M$, $j=1,\ldots,N$}$
    if ${\cal O}_{T}$ and ${\cal O}_{S}$ are causally separated. $\blacksquare$
This postulate requires invariance of the statistics of two consecutive selective measurements of the observables under interchange of their order. In fact, $\mathrm{tr}(\rho^{T}_{j}S_{i})\mathrm{tr}(\rho^{T}_{j})$ is the probability of measuring first $j$ for $T$ and next $i$ for $S$, whereas $\mathrm{tr}(\rho^{S}_{i}T_{j})\mathrm{tr}(\rho^{S}_{i})$ is the probability of measuring first $i$ for $S$ and next $i$ for $T$, in both cases the measured state being $\rho$. The temporal order of two measurements located in spatially separated regions has no physical meaning, and thus the statistics of the said measurements should be invariant under interchange of order.
3.3 NSC and RCC for Lüders and more sophisticated types of measurements
To continue the analysis started by Busch and Singh [BuSi98] and Busch [Bus99] (see also Section 2.2 of Beck’s thesis [Bec20] for developments and details) on the interplay among commutativity of effects, NSC, and RCC, we need some further general notions from quantum measurement theory. A special case of efficient measurement, which generalizes (finite discrete) projective measurements, is a Lüders measurement.
Definition 3.4:
[…]
Remark 3.5:
In the general case, directly from the polar decomposition theorem (see e.g. [Mor18]), any (Kraus) operator $K\in{\mathfrak{B}}({{\cal H}})$ satisfying $K^{\dagger}K=T$ (e.g., associated with an efficient measurement of the POVM $\{T,I-T\}$) can be written as $K=V\sqrt{T}$, where $V$ is a partial isometry with initial space $\mathrm{Ker}(V)^{\perp}=\mathrm{Ker}(\sqrt{T})^{\perp}=\mathrm{Ker}(T)^{\perp}=\overline{\mathrm{Ran}(\sqrt{T})}=\overline{\mathrm{Ran}(T)}\>.$ If $\mathrm{Ker}(T)$ is trivial, i.e., $\mathrm{Ran}(T)$ is dense, then $V$ is necessarily an isometry. However, even if it is an isometry, in general $V$ is not unitary unless $K$ is bijective. $\blacksquare$
We can now present the two fundamental results in which, in physical terms, locality formalized in terms of NSC and RCC is proved to imply commutativity of the effects of pairs of POVMs localized in causally separated regions.
Before stating these results, a comment is necessary. As clearly appears in the thorough analysis by Beck [Bec20] – see fig.3 of [Bec20] in particular – the two physical formulations of the locality principle, NSC and RCC, are not directly related to commutativity of the said effects. A more direct relation takes place between the effects of one observable and the Kraus operators of the other. In turn, this second type of commutativity implies commutativity of the effects. A more direct relation between effects and Kraus operators exists for Lüders measurements or similar measurement procedures.
A first general result due to Busch and Singh, which extends a previous result by Lüders [Lud51] and connects NSC to commutativity, reads as follows:
Proposition 3.6:
[…]
$T_{j},S]=0\quad\forall j\quad\Leftrightarrow\quad\mathrm{tr}\left(\rho S\right)=\mathrm{tr}\left(\rho_{L}^{T}S\right)\quad\mbox{for every state $\rho\in{\mathsf{S}}({{\cal H}})$,}$
where the measurement of $T$ is supposed to be non-selective and of Lüders type, holds in at least the following cases:
  • (i)
    $S$ has a discrete spectrum with eigenvalues that can be ordered in decreasing order, $T$ arbitrary;
  • (ii)
    $N=2$, so that $T_{2}=I-T_{1}$, and $S$ is arbitrary.
Proof.
This is Proposition 1 in [Bus99], which was established in [BuSi98]. ∎
A similar result can be proved referring to RCC, as announced in [Bus99] without proof. However, it holds in an even stronger version, as established in [GuNa01] (see the discussion in [Bec20]).
Proposition 3.7:
[…]
$T_{j},S_{i}]=0\quad\forall i,j\quad\Leftrightarrow\quad\mathrm{tr}\left(\rho^{T}_{L,j}S_{i}\right)\mathrm{tr}\left(\rho^{T}_{L,j}\right)=\mathrm{tr}\left(\rho^{S}_{L,i}T_{j}\right)\mathrm{tr}\left(\rho^{S}_{L,i}\right)\quad\forall i,j,\rho.$
More generally, the implication $\Leftarrow$ is valid for efficient measurement procedures where $K^{T}_{j}=K^{T\dagger}_{j}$ and $K^{S}_{i}=K^{S\dagger}_{i}$ for all $i,j$.
Proof.
Suppose that the identity on the left-hand side of the double implication is true. For Lüders measurements, $T_{j},S_{i}]=0$ (for all $i,j$) is equivalent to commutativity of the corresponding Kraus operators of one observable with the effects of the other, as immediately follows from the functional calculus of bounded selfadjoint operators (see, e.g., [Mor19]). Since these Kraus operators are in particular selfadjoint, the right-hand side of the double implication must be true as well, simply by applying the relevant definitions. The statement on the right-hand side of the double implication, for generic efficient measurements, is equivalent to $K^{T\dagger}_{j}S_{i}K^{T}_{j}=K^{S\dagger}_{i}T_{j}K^{S}_{i}$ (for every $i,j$). If the Kraus operators are also selfadjoint (as, in particular, for Lüders measurements), we can apply Theorem 2.10 in [Bec20] (arising from [GuNa01]), concluding that the identity on the left-hand side is true as well. ∎
When $T$ and $S$ are a pair of POVMs localized in causally separated regions, RCC and the above theorem imply that the effects must commute if the measurement process is efficient and the Kraus operators are selfadjoint, in particular for a Lüders measurement procedure. Similarly, if $S$ is itself an effect and the POVM $T$ and $\{S,I-S\}$ are localized in causally separated regions, NSC and Proposition 3.6 imply that the effects must commute if the measurements are of Lüders type. This substantial physical equivalence of no-signaling, consistency, and commutativity was already known for projective measurements [Mit98].
Remark 3.8:
[…]
$[\hat{\phi}[f],\hat{\phi}[f^{\prime}]]=0\qquad\mbox{if $supp(f)\subset{\cal O}$ and $supp(f^{\prime})\subset{\cal O}^{\prime}$, with ${\cal O},{\cal O}^{\prime}$ causally separated,}$ (5)
where the identity holds on a dense common invariant domain $\mathfrak{D}$$\subset{{\cal H}}$ of the *-algebra representation of the fields. The above identity is valid when NSC holds and:
  • (a)
    $\hat{\phi}[f],\hat{\phi}[f^{\prime}]$ are essentially self-adjoint operators in the field Hilbert space,
  • (b)
    they are supposed to be measured by projective measurements,
  • (c)
    the selfadjoint observables $F(\overline{\hat{\phi}[f]}),F^{\prime}(\overline{\hat{\phi}[f^{\prime}]})$ (i.e. their PVMs) are localized in $\cal O$} and $\cal O$$^{\prime}$ respectively for every choice of Borel real functions $F:{\mathbb{R}}\to{\mathbb{R}}$ and $F^{\prime}:{\mathbb{R}}\to{\mathbb{R}}$.
[…]
A final equivalence result, which extends Proposition 3.6 to generic non-efficient measurements of finite discrete POVMs and effects, exists due to Beck’s Theorem 3.12 [Bec20], but it uses a further, rather strong hypothesis.
Theorem 3.9 (Beck):
[…]
$K^{T}_{jk},S]=[K^{T\dagger}_{jk},S]=0\quad\forall j,k\quad\Leftrightarrow\quad\mathrm{tr}\left(\rho S\right)=\mathrm{tr}\left(\rho^{T}S\right)\quad\mbox{and}\quad\mathrm{tr}\left(\rho S^{2}\right)=\mathrm{tr}\left(\rho^{T}S^{2}\right)\quad\forall\rho\>,$
where we consider a generic non-selective measurement of $T$ according to (4).
Remark 3.10:
$K^{T}_{jk},S]=[K^{T\dagger}_{jk},S]=0$ immediately implies $T_{j},S]=0$ from the latter identity in (3). $\blacksquare$
4 Commutativity requirement and relativistic locality
We are now in a position to study the relation between Propositions 3.6, 3.7 and the validity of the commutativity hypothesis (4) in Halvorson-Clifton’s theorem, also for relativistic spatial localization observables and, in this case, also comparing effects with different tensorial indices $v$. First, we prove that, in principle, these theorems – where NSC/RCC and Lüders measurements play a crucial role – imply commutativity for the effects of an unsharp localization system associated with causally separated regions, provided some apparently natural hypotheses are fulfilled. Next, we prove that causal separation hypotheses are, however, never satisfied simply in view of the very physical nature of a particle and its propensity to localize in a unique place. Hence the commutativity assumption does not arise by this route. In this case, failure of commutativity is not in contradiction with locality principles such as NSC and RCC.
4.1 NSC, RCC and commutativity of effects
Busch’s crucial result about condition (4) in the Halvorson-Clifton theorem can now be stated precisely, also in a more general version which takes Proposition 3.7 into account.
Theorem 4.1:
Consider two regions $\Delta,\Delta^{\prime}\in{\cal R}$ for either a relativistic spatial localization observable or an unsharp localization system $({{\cal H}},{\cal R},A,U)$ (in this case the specifications $v,v^{\prime}$ below are not necessary), and assume there exist open regions of spacetime ${\cal O}_{\Delta}\supset\Delta$ and ${\cal O}_{\Delta^{\prime}}\supset\Delta^{\prime}$ such that
  • (i)
    the POVM $\{A^{v}(\Delta),I-A^{v}(\Delta)\}$ is localized in ${\cal O}_{\Delta}$;
  • (ii)
    the POVM $\{A^{v^{\prime}}(\Delta^{\prime}),I-A^{v^{\prime}}(\Delta^{\prime})\}$ is localized in ${\cal O}_{\Delta^{\prime}}$.
If moreover,
  • (a)
    measurements of the said POVMs are performed using Lüders procedures,
  • (b)
    ${\cal O}_{\Delta}$ and ${\cal O}_{\Delta^{\prime}}$ are causally separated,
  • (c)
    NSC or RCC holds (in the latter case the measurement procedures in (a) can be weakened to efficient measurements with selfadjoint Kraus operators),
then $A^{v}(\Delta),A^{v^{\prime}}(\Delta^{\prime})]=0.$
Proof.
This immediately follows from case (ii) of Proposition 3.6, with the POVM $T$ described by $\{A^{v}(\Delta),I-A^{v}(\Delta)\}$ and the effect $S$ given by $A^{v^{\prime}}(\Delta^{\prime})$. Alternatively, it immediately follows from Proposition 3.7, with $S$ interpreted as the POVM $\{A^{v}(\Delta^{\prime}),I-A^{v^{\prime}}(\Delta^{\prime})\}$. ∎
Remark 4.2:
An alternative result may be proved by taking advantage of Theorem 3.9 and Remark 3.10. Compared with Theorem 4.1, one could now consider much more general measurement procedures: non-efficient and given in terms of non-selfadjoint operators. Similarly to Theorem 4.1, hypotheses (i), (ii), (a), (b) should be assumed, as well as (c) in the case of NSC. However, to exploit the implication $\Leftarrow$ of Theorem 3.9 together with Remark 3.10, one should assume, for every $\rho\in{\mathsf{S}}({{\cal H}})$ (where we omit the possible specifications $v,v^{\prime}$),
$\mathrm{tr}\left(\rho A(\Delta^{\prime})^{2}\right)=\mathrm{tr}\left(\rho^{T}A(\Delta^{\prime})^{2}\right)\quad\mbox{with $T$ given by $\{A(\Delta),I-A(\Delta)\}$}$
which, in principle, might again be justified as a consequence of NSC. However, in the identity above, $A(\Delta^{\prime})^{2}$ may not be viewed as an elementary effect, but rather as a selfadjoint observable in its own right (with its own PVM), instead of encapsulating it in an elementary POVM $\{A(\Delta^{\prime}),I-A(\Delta^{\prime})\}$. In particular, according to Definition 3.1, we should also assume that there is a further physical procedure to measure the PVM of the selfadjoint operator $A(\Delta^{\prime})^{2}$ and that this PVM is localized in a region causally separated from $\cal O_{\Delta}$}. In this case, using the standard elementary formulation of quantum theory (see, e.g., [Mor19]), $\mathrm{tr}\left(\rho A(\Delta)^{2}\right)$ is interpreted as the expectation value $\langle A(\Delta^{\prime})^{2}\rangle_{\rho}$ of the selfadjoint observable $A(\Delta)^{2}$ in the state $\rho$. In the AHK framework, adopting this additional hypothesis also seems fairly natural, but the use of the AHK setting is exactly what we are trying to avoid. An alternative possibility is to introduce another elementary POVM $\{A(\Delta^{\prime})^{2},I-A(\Delta^{\prime})^{2}\}$ and assume that it is localized in a region causally separated from ${\cal O}_{\Delta}$. From a general perspective, the fact that the mathematical expression of the no-signaling condition may hold for an effect $S$: $\mathrm{tr}\left(\rho S\right)=\mathrm{tr}\left(\rho^{T}S\right)$ for every $\rho\in{\mathsf{S}}({{\cal H}})$, but not for $S^{2}$: $\mathrm{tr}\left(\rho S^{2}\right)\neq\mathrm{tr}\left(\rho^{T}S^{2}\right)$ for some $\rho\in{\mathsf{S}}({{\cal H}})$, is a concrete possibility. A counterexample due to Heinosaari and Wolf [HeWo10] is discussed by Beck in Section 2.5 of [Bec20]. $\blacksquare$
In view of Theorem 4.1, it is then straightforward to see that requirement (4) in Halvorson-Clifton’s theorem 2.5 holds under some further reasonable physical assumptions:
  • (i)
    all measurements of the POVMs $\{A(\Delta),I-A(\Delta)\}$, for $\Delta\in{\cal R}$, are performed using the Lüders procedure assuming NSC – or with efficient selfadjoint Kraus operators and assuming RCC;
  • (ii)
    for any $\Delta\in{\cal R}$, $\{A(\Delta),I-A(\Delta)\}$ can be localized in open regions ${\cal O}^{\epsilon}_{\Delta}\supset\Delta$ that shrink around $\Delta$ as $\epsilon\to 0^{+}$: ${\cal O}^{\epsilon}_{\Delta}$ is the causal completion of a covering of $\Delta$ by open balls of radius $\epsilon>0$ within the $3$-plane $\Sigma\ni\Delta$.
[…]
4.2 Commutativity and Relativistic Spatial Localization Observables
We want to study the validity of hypothesis (4) in the case where the set of effects $A^{v}(\Delta)$ is organized into, or extends to, families of POVMs forming a relativistic spatial localization observable according to Definition 2.2. However, this structure may be considerably weakened by using an extension of the notion of unsharp localization system where (1) unbounded sets $\Delta$ are allowed and (2) $\Sigma\setminus\Delta\in{\cal R}$ if $\Delta\in{\cal R}$ and $\Delta\subset\Sigma$. Yet, requirements (c) and (d) in Definition 2.2 do not play any role here.
Proposition 4.3:
Let $({{\cal H}},{\cal R},A,U)$ be a relativistic spatial localization observable, let $\Sigma$ a rest space of a reference frame, and suppose that LDP holds. If, for $\Delta\in{\cal R}$ with $\Delta\subset\Sigma$, the elementary POVM $\{A^{v}(\Delta),I-A^{v}(\Delta)\}$ is localized in the open set ${\cal O}\subset{\mathbb{M}}$, then ${\cal O}\supset\Sigma$ necessarily.
Proof.
Using the definition of POVM, $\{A^{v}(\Delta),I-A^{v}(\Delta)\}=\{I-A^{v}(\Delta^{c}),A^{v}(\Delta^{c})\}$ where $\Delta^{c}:=\Sigma\setminus\Delta\in{\cal R}$ due to the very definition of relativistic spatial localization observable. At this point, since $\{A^{v}(\Delta^{c}),I-A^{v}(\Delta^{c})\}$ is localized in $\cal O$}, LDP implies that $\Delta^{c}\subset{\cal O}$. In summary, ${\cal O}\supset\Delta\cup\Delta^{c}=\Sigma$. ∎
Corollary 4.4:
Let $({{\cal H}},{\cal R},A,U)$ be a relativistic spatial localization observable, suppose that LDP holds, and consider $\Delta,\Delta^{\prime}\in{\cal R}$. If the correspondingly associated elementary POVMs $\{A^{v}(\Delta),I-A^{v}(\Delta)\}$ and $\{A^{v^{\prime}}(\Delta^{\prime}),I-A^{v^{\prime}}(\Delta^{\prime})\}$ are respectively localized in ${\cal O}_{\Delta}$ and ${\cal O}_{\Delta^{\prime}}$, then these sets cannot be causally separated (even if $\Delta$ and $\Delta^{\prime}$ are causally separated).
Proof.
As ${\cal O}_{\Delta}\supset\Sigma$ and the latter is a spacelike Cauchy surface, $J^{+}({\cal O}_{\Delta^{\prime}})\cup J^{-}({\cal O}_{\Delta^{\prime}})$ must meet $\Sigma$ – and thus ${\cal O}_{\Delta}$ – somewhere. ∎
Consequently, $\{A^{v}(\Delta),I-A^{v}(\Delta)\}$ and $\{A^{v^{\prime}}(\Delta^{\prime}),I-A^{v^{\prime}}(\Delta^{\prime})\}$ do not satisfy the basic hypothesis (b) of Theorem 4.1. We conclude that commutativity, i.e., the natural generalization of hypothesis (4) in Halvorson-Clifton’s theorem, is not physically mandatory for relativistic spatial localization observables and for unsharp localization systems that are extendable to the former type of structure. This result is actually quite general. Even establishing a more general version of Theorem 4.1, also taking Remark 4.2 into account and considering measurement procedures much more general than Lüders(-like) ones, would not remove the obstruction found to a proof of commutativity from locality. Corollary 4.4 shows that, if a relativistic spatial localization observable exists, the region where elementary observables are localized is the whole physical space at the corresponding instant of time, even if the observables refer to bounded detection regions. Two localization regions, unlike the detection regions, are never causally separated, and this does not permit one to trigger locality arguments relying upon NSC or RCC. The apparently only alternative route is to abandon from the outset the physically natural principle LDP, which played a crucial role in the proof of Proposition 3.7.
In summary, what we have established with Proposition 4.3 and Corollary 4.4 has a clear physical meaning. We are considering systems which localize in a unique region, namely at a single point of space in ideal detection procedures, and not in several regions simultaneously: in other words, we are considering a particle. A particle has the propensity to localize in a single place, unlike extended systems [^11] Since localization arises from a physical interaction, it is nevertheless questionable whether localization in a region smaller than the particle’s Compton wavelength is physically meaningful, in view of the possibility of pair creation.. We are simply exploring the consequences in measurement theory of this fundamental propensity characteristic of particle nature. Using a POVM on $\Sigma$ of a relativistic spatial localization observable means that we have filled the entire rest space $\Sigma$ of an inertial reference frame with infinitely many detectors. We turn them on at the instant $t$ – which defines $\Sigma$ – and turn them off immediately afterwards. At that instant, the the particle must necessarily be detected in a unique detector. If we focus our attention on an arbitrarily chosen region $\Delta\subset\Sigma$, the system may or may not be found there at time $t$. In the first case, however, we are sure that it is simultaneously detected nowhere in $\Delta^{c}$; in the second case, we are sure that it is simultaneously detected somewhere in $\Delta^{c}$. By observing what happens in $\Delta$, we therefore have information about what is happening in the whole of $\Sigma$ at the instant $t$. In this sense, it seems natural that $\{A^{v}(\Delta),I-A^{v}(\Delta)\}$ should be thought of as localized everywhere in $\Sigma$, and there is no contradiction with locality because, as said above, we are using the propensity of the considered quantum system, namely a particle, to localize somewhere and in a unique place if its position is measured. If we admit that the system is allowed to localize “simultaneously” in several places, i.e., in causally separated regions, we are no longer using the notion of particle and the notion of relativistic spatial localization observable is not appropriate for such a system. Whether or not particles exist, in the said sense, is simply a matter for experimental investigation. In this work, we assume that they do exist. In a forthcoming paper we shall more strongly assume that elementary particles in Wigner’s sense are of this sort, at least massive scalar bosons.
This analysis can also be applied to the case where the regions $\Delta$ and $\Delta^{\prime}$ belong to different rest spaces of different reference frames, provided these regions are causally separated. This is a consequence of the fact that [DRMo24] (at least some) relativistic spatial localization observables can be extended to the family of generic smooth spacelike Cauchy surfaces. Furthermore, for spacelike separated (sufficiently regular) spacelike subsets $\Delta$ and $\Delta^{\prime}$, there is a spacelike Cauchy surface $S$ which contains both [BeSa06]. We can use $S$ as our notion of rest space in which to detect the particle, and we can recast the previous argument, this time assuming that theis generalized rest space $S$ is filled with detectors. The result is even more generalizable by taking $\Delta$ and $\Delta^{\prime}$ of minimal regularity (Lipschitz achronal Borel sets) and exploiting a maximal achronal set $S$ which includes both [CDM26]. Finally, even in the situation where we switch on only two detectors in causally separated regions $\Delta$ and $\Delta^{\prime}$ of a given rest space $\Sigma$, the above reasoning could hold in a weaker version. If the particle is found in $\Delta$, it cannot be found in $\Delta^{\prime}$ simply in view of its propensity to localize in a unique place. Unlike the previous case, however, there is now the possibility that the particle localizes in none of the switched-on detectors, since they do not exhaust all possibilities. However, this situation is more delicate to analyze, and we do not address it here.
As a consequence of the discussion above, in an AHK perspective, we expect that every $A^{v}(\Delta)$ of a relativistic spatial localization observable, if any, does not belong to a local algebra of observables ${\mathfrak{A}}({\cal O})$ with $\cal O$} arbitrarily narrowed around $\Delta$. Otherwise, commutativity should hold for sufficiently causally separated $\Delta$ and $\Delta^{\prime}$. This $\cal O$} has to be taken as a neighborhood of the whole $\Sigma$ and cannot be restricted further.
It is eventually important to emphasize that we are not claiming that hypothesis (4) of H-C Theorem 2.5 cannot hold in general. Rather, we argue that, from a physical standpoint, commutativity of effects of localization observables, if any, is not a consequence of NSC or RCC (through Propositions 3.6,3.7, and 4.1 in particular). Indeed, there exist in the literature relativistic spatial localization observables concerning a particle-antiparticle fermion [Cas17] and described by families of PVMs so that, for every $\Delta,\Delta^{\prime}\subset\Sigma$, it must be $A(\Delta),A(\Delta^{\prime})]=0$ – and more precisely $A(\Delta)A(\Delta^{\prime})=A(\Delta^{\prime})A(\Delta)=A(\Delta\cap\Delta^{\prime})=0$ if $\Delta\cap\Delta^{\prime}=\varnothing$. This is true on account of the very definition of PVM and not as a consequence of Busch’s theorem. For these families of PVMs, hypothesis (4) also holds as a consequence of standard properties of orthogonal projectors and CC (see a proof in Section A). For these relativistic spatial localization observables concerning a fermion [Cas17], however, the conclusion of Halvorson-Clifton’s Theorem 2.5 does not hold, because in general $A(\Delta)\neq 0$. The reason is that the energy-boundedness requirement (3) in Theorem 2.5 is not satisfied. Furthermore, upon restriction to the particle space, namely the positive-energy sector, commutativity ceases to hold, and the observables are therefore described by POVMs.
Nevertheless, if one insists on maintaining also the energy-boundedness requirement together with additivity and translation covariance, then commutativity cannot hold simply because of the Halvorson-Clifton result. However, for a quantum particle, understood as above, this fact does not seem to pose a problem.
5 Possible commutativity recovery via local conditional POVMs
In this section, we propose a different viewpoint on localization of quantum systems, where measurements are explicitly performed in finite regions of space, namely laboratories. By construction, the observables we are going to define should in principle be causally independent when associated with causally separated spatial regions. We focus on probabilities of detecting the system in one of the detectors of a laboratory, when it is already known that the system is found in the laboratory. Hence the physical interpretation of the procedure we are about to introduce involves conditional probabilities.
5.1 Localization in laboratories
Realistic localization experiments in $\mathbb{M}$} are made in laboratories, which are spatially finite regions. We shall also assume that they are causally complete in their spacetime description: everything that may happen in such a spacetime region should be determined by physical actions performed therein, at least at a macroscopic level. Let $\Sigma$ be a flat $3$-dimensional plane. We define a laboratory $L(\Delta_{0})\subset{\mathbb{M}}$, with space given by a bounded set $\Delta_{0}\in\mathscr{B}(\Sigma)$, as the causal completion [^12] We warn the reader that this orthogonality relation is not the one used in [CDM26], where achronal separation was instead considered to describe the orthomodular lattice of the so-called causal logic of Minkowski spacetime. defined as in (1)
$L(\Delta_{0}):=(\Delta_{0}^{\perp\!\!\!\!\perp})^{\perp\!\!\!\!\perp}\>.$ (6)
In particular, if $\Delta_{0}$ is open in the relative topology, the resulting open set $L(\Delta_{0})$, equipped with the Lorentzian manifold structure induced by $\mathbb{M}$}, is a globally hyperbolic spacetime in its own right, with $\Delta_{0}$ as a possible smooth spacelike Cauchy surface. $L(\Delta_{0})$ also contains other curved smooth spacelike Cauchy surfaces, though we shall focus only on the bounded flat 3-space $\Delta_{0}$ of the laboratory $L(\Delta_{0}).$
Given a laboratory $L(\Delta_{0})$, we want to localize a quantum system in the subregions $\Delta\subset\Delta_{0}$ of its flat bounded portion of rest space $\Delta_{0}$, even if the system can generally be detected in the whole spacetime. For the moment, we once more assume that the entire space is filled with detectors represented by a relativistic spatial localization observable $A$. In Section 5.4 we shall try to get rid of this unrealistic perspective. The idea is to define a POVM referring only to the laboratory, starting from a notion of localization $A$ defined on the whole space $\Sigma$. This “restricted” POVM should describe in which subregion $\Delta\in\mathscr{B}(\Delta_{0})$ of the laboratory the system is detected if the system is detected in $\Delta_{0}$. Notice that the POVM we are looking for should be normalized on $\Delta_{0}$ and not on the whole $\Sigma$. It concerns the fraction of events in $\Delta$ over the events in $\Delta_{0}$ (and not over those in the whole $\Sigma$). If we denote the effects of this POVM by $B_{\Delta_{0}}(\Delta)$, then $tr(\rho B_{\Delta_{0}}(\Delta))$ is the probability of finding the system in $\Delta$ if it is found in $\Delta_{0}$. In particular, $tr(\rho B_{\Delta_{0}}(\Delta_{0}))=1$ because it is the probability of finding the system in $\Delta_{0}$ if it is found in $\Delta_{0}$!
5.2 Localization in a laboratory out of a relativistic spatial localization observable
Suppose we have at our disposal a full relativistic spatial localization observable denoted by $A^{v}:{\cal R}\to{\mathfrak{B}}({{\cal H}})$ or just a POVM $A$ on a specific spacelike $3$-plane $\Sigma$. Take $\Delta_{0}\in\mathscr{B}(\Sigma)$. We intend to re-normalize $A$ on this $\Delta_{0}$. The idea is to consider families of operators of the form
$B_{\Delta_{0}}(\Delta)=\frac{1}{\sqrt{A(\Delta_{0})}}A(\Delta)\frac{1}{\sqrt{A(\Delta_{0})}}\qquad\mbox{where $\Delta\in\mathscr{B}(\Delta_{0})$.}$ (7)
That is because, if these operators can be constructed, they define a normalized POVM on $\Delta_{0}$. However, some technical issues arise immediately in connection with the proposed definition above, in particular the existence of the inverse operators written there. Let us examine the less obvious ingredients of the formula above. From functional calculus (and selfadjointness) it is easy to prove that $Ker(\sqrt{A(\Delta_{0})})=Ker(A(\Delta_{0}))$ simply because ${\mathfrak{B}}({{\cal H}})\ni A(\Delta_{0})\geq 0$. Therefore, if $Ker(A(\Delta_{0}))=\{0\}$, then a selfadjoint operator $A(\Delta_{0})^{-1/2}:Ran(\sqrt{A(\Delta_{0})})\to{{\cal H}}$ turns out to be well defined. Notice that the domain is dense because $\overline{Ran(\sqrt{A(\Delta_{0})})}=Ker(\sqrt{A(\Delta_{0})})^{\perp}={{\cal H}}$. At this point, an important technical result concerning the condition $Ker(A(\Delta_{0}))=\{0\}$ holds, which arises from even stronger results due to Castrigiano [^13] I am grateful to Castrigiano for some clarifications about this issue. [Cas17] extending previous results by Hegerfeldt. In the next proposition, the whole structure of a relativistic spatial localization observable $({{\cal H}},A,{\cal R},U)$ is not necessary.
Proposition 5.1:
[…]
  • (1)
    Referring to the above Minkowskian reference frame $u$, the square-mass operator
    $M^{2}:=P_{0}^{2}-\sum_{k=1}^{3}P_{k}^{2}$
    of the strongly-continuous unitary representation $U$ of $\widetilde{ISO(1,3)_{+}}$ (defined in the Gårding domain [^14] It is essentially selfadjoint theoreon as every central element of the universal enveloping algebra of the Lie algebra of $\widetilde{ISO(1,3)_{+}}$ represented in terms of generators of $U$. This condition does not depend on the used Minkowskian reference frame.) is non-negative.
  • (2)
    Energy is positive: the selfadjoint generator $H^{u}=-P_{0}=P^{0}$ of time evolution along $u$ through $U$ (i.e. past-directed temporal translation along $u$) has non-negative spectrum.
  • (3)
    CT is satisfied for the rest spaces $\Sigma_{u}$ normal to $u$.
[…]
$Ker(A(\Delta_{0}))\neq\{0\}\quad\Rightarrow\quad\overline{\Sigma_{u}\setminus\Delta_{0}}=\Sigma_{u}\>.$
If $({{\cal H}},A,{\cal R},U)$ is a relativistic spatial localization observable which satisfies CC and conditions (1) and (2) (which do not depend on the reference frame), then the thesis is true for every rest space $\Sigma$ and every $A^{v}(\Delta)$ with $\Delta\subset\Sigma$.
Proof.
Saying that $Ker(A(\Delta_{0}))\neq\{0\}$ is equivalent to stating that $A(\Delta_{0})\psi=0$ for some $\psi\in{{\cal H}}\setminus\{0\}$ which, in turn, is equivalent to $A(\Delta_{0}^{c})\psi=\psi$, where $\Delta_{0}^{c}:=\Sigma_{u}\setminus\Delta_{0}$. The condition $A(\Delta_{0}^{c})\psi=\psi$ for $\psi\neq 0$ implies, in fact, that $\Delta_{0}^{c}$ is essentially dense in $\Sigma_{u}$: $\overline{\Delta^{c}_{0}\setminus N}=\Sigma_{u}$ for every zero-Lebesgue-measure set $N\subset\Sigma_{u}$ as a consequence of (8) Theorem in [Cas17], also taking the comment under (5) Definition therein into account. Using $N=\varnothing$, we obtain $\overline{\Sigma_{u}\setminus\Delta_{0}}=\Sigma_{u}$, which is the thesis. The last statement is a trivial consequence of what has been proved and of the definition of relativistic spatial localization observable. ∎
Remark 5.2:
This type of result agrees with the idea that a relativistic quantum system cannot be sharply localized – with probability $1$ – in a non-dense spatial region, in particular a bounded one: the probability of finding it arbitrarily far away in space is always positive, even if it tends to vanish, when time evolution is causal (and energy and the square-mass operator are positive, as expected for relativistic systems). An important issue is whether, conversely, localized probability distributions can be approximated arbitrarily well by admissible states of the said type. A general discussion of these issues, as well as concrete examples for single fermions, appears in [Cas17]. For single bosons, some results are presented in [Mor23, Cas24]. $\blacksquare$
We are now in a position to give a rigorous technical interpretation of the right-hand side of (7), even in a slightly more general version.
Lemma 5.3:
Let $A:\mathscr{B}(\Sigma)\to{\mathfrak{B}}({{\cal H}})$ be a POVM, where $\Sigma\subset{\mathbb{M}}$ is a spacelike $3$-plane. Choose
  • (1)
    $\Delta_{0}\in\mathscr{B}(\Sigma)$ such that $Ker(A(\Delta_{0}))=\{0\}$ [^15] This is in particular true if $\Delta_{0}$ is bounded with $Int(\Delta_{0})\neq\varnothing$, and (1),(2),(3) of Proposition 5.1 hold.;
  • (2)
    an isometry $V:{{\cal H}}\to{{\cal H}}$ with final space $Ran(V)(=\overline{Ran(V)})=:{{\cal H}}_{V}$.
For every $\Delta\in\mathscr{B}(\Delta_{0})$ there is a unique operator $B^{V}_{\Delta_{0}}(\Delta)\in{\mathfrak{B}}({{\cal H}}_{V})$ such that
$\left\langle\frac{1}{\sqrt{A(\Delta_{0})}}V^{\dagger}\psi\left|A(\Delta)\frac{1}{\sqrt{A(\Delta_{0})}}V^{\dagger}\phi\right.\right\rangle=\langle\psi|B^{V}_{\Delta_{0}}(\Delta)\phi\rangle\qquad\mbox{if $\psi,\phi\in V(Ran(\sqrt{A(\Delta_{0})}))$.}$ (8)
Furthermore, $B^{V}_{\Delta_{0}}(\Delta)$ is an effect with $B^{V}_{\Delta_{0}}(\Delta_{0})=VV^{\dagger}$, the orthogonal projector onto the final space of $V$.
Proof.
By hypothesis
$0\leq A(\Delta)+A(\Delta_{0}\setminus\Delta)=A(\Delta_{0})\leq I$
and thus $A(\Delta)\leq A(\Delta_{0})$ (because $A(\Delta_{0}\setminus\Delta)\geq 0$). From $A(\Delta)\leq A(\Delta_{0})$ and using the standard functional calculus [Mor18], we have that
[…]
$\langle V^{\dagger}\psi|V^{\dagger}\psi\rangle=\langle\psi|\psi\rangle\quad\mbox{if $\psi\in V(Ran(\sqrt{A(\Delta_{0})}))$.}$
where we used the fact that $VV^{\dagger}$ is the orthogonal projector onto ${{\cal H}}_{V}=Ran(V)=\overline{Ran(V)}$ since $V:{{\cal H}}\to{{\cal H}}$ is an isometry. A straightforward use of the Cauchy-Schwarz inequality for the positive sesquilinear form inside the absolute value below, together with the previous inequality, proves that
[…]
Since $Ran(\sqrt{A(\Delta_{0})})$ is dense, from Riesz’ lemma we conclude that there exists a unique operator $B^{V}_{\Delta_{0}}(\Delta)\in{\mathfrak{B}}({{\cal H}}_{V})$ such that (8) holds:
$\left\langle\frac{1}{\sqrt{A(\Delta_{0})}}V^{\dagger}\psi\left|A(\Delta)\frac{1}{\sqrt{A(\Delta_{0})}}V^{\dagger}\phi\right.\right\rangle=\langle\psi|B^{V}_{\Delta_{0}}(\Delta)\phi\rangle\quad\mbox{if $\psi,\phi\in V(Ran(\sqrt{A(\Delta_{0})}))$,}$
where the closed subspace ${{\cal H}}_{V}:=Ran(V)=\overline{V(Ran(\sqrt{A(\Delta_{0})}))}\subset{{\cal H}}$ is viewed as a Hilbert space in its own right. Furthermore, by continuity (in both formulas $I$ is the identity operator in ${{\cal H}}_{V}$)
$0\leq B^{V}_{\Delta_{0}}(\Delta)\leq I\qquad\mbox{and}\qquad B^{V}_{\Delta_{0}}(\Delta_{0})=VV^{\dagger}$ (9)
are valid. ∎
Remark 5.4:
If, in addition to the hypotheses above, we also know that $A(\Delta_{0})\geq cI$ for some $c>0$, we can conclude that $Ran(A(\Delta_{0}))={{\cal H}}$ and $\frac{1}{\sqrt{A(\Delta_{0})}}\in{\mathfrak{B}}({{\cal H}})$. In this situation
$B^{V}_{\Delta_{0}}(\Delta)=V\frac{1}{\sqrt{A(\Delta_{0})}}A(\Delta)\frac{1}{\sqrt{A(\Delta_{0})}}V^{\dagger}$ (10)
is an equivalent version of (8). $\blacksquare$
The effects $B^{V}_{\Delta_{0}}(\Delta)$ are our candidates for defining a notion of conditional spatial localization in the rest space $\Delta_{0}$ of a laboratory. To avoid issues with normalization of the POVM we want to define, it is convenient to assume that $VV^{\dagger}=I$, which is equivalent to requiring that $V$ be unitary. The final result can now be stated and proved while considering the general case.
Proposition 5.5:
Let $A$ be a relativistic spatial localization observable and $V:{{\cal H}}\to{{\cal H}}$ a unitary operator. If $\Delta_{0}\in\mathscr{B}(\Sigma)$, with $\Sigma\subset{\mathbb{M}}$ a spacelike $3$-plane, satisfies $|\Delta_{0}|>0$ and $Ker(A^{v}(\Delta_{0}))=\{0\}$ for every $v\in{\mathsf{S}}$. Then
  • (a)
    $\mathscr{B}(\Delta_{0})\ni\Delta\mapsto B^{v,V}_{\Delta_{0}}(\Delta)\in{\mathfrak{B}}({{\cal H}})$ with $B^{v,V}_{\Delta_{0}}(\Delta)$ defined in (8) for $A^{v}$ in place of $A$, is a (normalized) POVM as well, which is absolutely continuous with respect to the Lebesgue measure on $\Sigma$;
  • (b)
    If the unitary representation of the universal covering of the proper orthochronous Poincaré group used in the definition of $A$ is $U:\widetilde{\cal P}_{+}\to{\mathfrak{B}}({{\cal H}})$, the covariance relation holds
    $U_{g}B^{v,V}_{\Delta_{0}}(\Delta)U_{g}^{-1}=B^{gv,U_{g}VU_{g}^{-1}}_{g\Delta_{0}}(g\Delta)\quad\mbox{for every $g\in\widetilde{\cal P}_{+}$, $\Delta\in\mathscr{B}(\Delta_{0})$.}$
Proof.
(a) We omit the tensorial index $v$ for notational simplicity. Taking (9) into account, the thesis holds if, for every $\psi\in{{\cal H}}$, the map
$\mathscr{B}(\Delta_{0})\ni\Delta\mapsto\langle\psi|B^{V}_{\Delta_{0}}(\Delta)\psi\rangle=:\mu_{\psi}(\Delta)\in[0,+\infty)$
[…]
(b) It easily follows from (d) of Definition 2.2, equation (8), and the spectral calculus of the spectral functions of the selfadjoint operator $A^{v}(\Delta_{0})$, using the fact that (d) of Definition 2.2 implies: $Uf(A^{v}(\Delta))U^{-1}=f(UA^{v}(\Delta)U^{-1})$ for every Borel measurable $f:{\mathbb{R}}\to{\mathbb{C}}$ and every unitary $U\in{\mathfrak{B}}({{\cal H}})$ [Mor19]. ∎
5.3 $B_{\Delta_{0}}(\Delta)$ as conditional POVM: the role of the gentle measurement lemma
Suppose that the initial state is $\rho$ and that we are measuring a scalar relativistic spatial localization observable $A$. For simplicity, we assume that $A$ is scalar, since what follows has an immediate generalization. As said previously, we are interested in the probability of finding the system – prepared in the state $\rho$ – in the Borel subsets $\Delta\subset\Delta_{0}$ when we know that the system is found in $\Delta_{0}$. Here $\Delta_{0}\subset\Sigma$ has finite Lebesgue measure and, to this end, we may assume that $\Delta_{0}$ is bounded. We also explicitly suppose that $Ker(A(\Delta_{0}))=\{0\}$, in order to define the new POVM $B^{V}_{\Delta_{0}}:\mathscr{B}(\Delta_{0})\to{{\cal H}}$ upon a choice of the unitary $V$.
To prove that $B^{V}_{\Delta_{0}}$ is a good candidate for representing this type of conditional probability, for some choice of $V$ at least, we adopt a heuristic approach. If we know that the system is found in $\Delta_{0}$, assuming that $A$ was measured with a measurement procedure described by Kraus operators as in Remark 3.5, we expect that the state has a form
${\rho}^{V}_{\Delta_{0}}:=\frac{V\sqrt{A(\Delta_{0})}{\rho}\sqrt{A(\Delta_{0})}V^{\dagger}}{tr(A(\Delta_{0})\tilde{\rho})}\in{\mathsf{S}}({{\cal H}})\>,$ (11)
for some other state ${{\rho}}\in{\mathsf{S}}({{\cal H}})$ and for some partial isometry $V$, which we explicitly suppose to be unitary and which, for the moment, is arbitrary. Notice that $tr(A(\Delta_{0}){\rho})\neq 0$ as a consequence of our hypothesis $Ker(A(\Delta_{0}))=\{0\}$ and ${\rho}\in{\mathsf{S}}({{\cal H}})$. At this point, a straightforward computation based on the cyclic property of the trace and (11) yields, for every $\rho\in{\mathsf{S}}({{\cal H}})$,
$\frac{tr(\rho A(\Delta))}{tr(\rho A(\Delta_{0}))}=tr\left(\rho^{V}_{\Delta_{0}}B^{V}_{\Delta_{0}}(\Delta)\right)\>.$ (12)
We have therefore found that the probability of localizing the system in $\Delta\subset\Delta_{0}$ – if we know that it was found in the laboratory rest space $\Delta_{0}\subset\Sigma$ – is precisely defined by the POVM $B^{V}_{\Delta_{0}}:\mathscr{B}(\Delta_{0})\to{{\cal H}}$. This heuristic reasoning, however, has two related shortcomings, since we assumed that the initially prepared state was $\rho$ and not $\rho^{V}_{\Delta_{0}}$, which in the discussion above plays only an intermediate role. Furthermore, the above heuristic interpretation refers to two measurements: one which transforms $\rho$} into ${\rho}^{V}_{\Delta_{0}}$ and a subsequent one to detect the system in some $\Delta\subset\Delta_{0}$. Rather, we want to study a situation in which only one measurement takes place on $\rho$: we only consider the events occurring in $\Delta_{0}$, disregarding those outside $\Delta_{0}$, and, without performing further measurements, we focus on the subset of such events occurring in the sets $\Delta\subset\Delta_{0}$. To address these issues, taking advantage of a suitable version of the so-called gentle measurement lemma, we are about to show that we can in fact replace ${\rho}^{V}_{\Delta_{0}}$ with $\rho$ in (12) provided the probability of finding the system in $\Delta_{0}$ for the initial state $\rho$ is close to $1$. All this is true provided $V=I$. (We shall come back to the case $V\neq I$ at the end of the section.) We use the original version of the result due to A. Winter [Win99]. The name “gentle measurement lemma” is more modern and nowadays also refers to an even tighter estimate in the finite-dimensional case. Winter’s proof of the following proposition was also carried out in the finite-dimensional case but, as is well known and as is clear from the proof itself [Win99], finite dimensionality plays no essential role.
Theorem 5.6 (Gentle measurement lemma):
[…]
Proof.
[…]
It is easy to see that the statement is generally false if one replaces $\sqrt{T}$ with $V\sqrt{A}$ for an effect $A$ and a unitary map $V$ which does not commute with $\rho$.
As a consequence of the above result, we assume for now that $V=I$ and, in applying the above theorem, we henceforth use the notation for this special case
$B^{v}_{\Delta_{0}}(\Delta):=B^{v,I}_{\Delta_{0}}(\Delta)\qquad\mbox{and, for the scalar case}\qquad B_{\Delta_{0}}(\Delta):=B^{I}_{\Delta_{0}}(\Delta)\>.$ (14)
Proposition 5.7:
[…]
If $\rho\in{\mathsf{S}}({{\cal H}})$ is such that the probability of finding the system in $\Delta_{0}$ satisfies
$tr(\rho A^{v}(\Delta_{0}))\geq 1-\delta>0\>,$
then
$\left|tr(\rho B^{v}_{\Delta_{0}}(\Delta))-\frac{tr(\rho A^{v}(\Delta))}{tr(\rho A^{v}(\Delta_{0}))}\right|\leq 2\sqrt{\delta}+\delta\qquad\mbox{for every measurable $\Delta\subset\Delta_{0}$.}$ (15)
More generally,
[…]
Proof.
[…]
We conclude that, for states whose probability of finding the system in $\Delta_{0}$ is close to $1$, $B_{\Delta_{0}}(\Delta)$ measures the conditional probability of finding the system in $\Delta$ if we know that it was found in $\Delta_{0}$. Statistically speaking,
$tr\left({\rho}B_{\Delta_{0}}(\Delta)\right)\simeq\frac{tr(\rho A(\Delta))}{tr(\rho A(\Delta_{0}))}\hskip 18.49988pt\mbox{where $\Delta\subset\Delta_{0}\qquad\mbox{and}\qquad\Delta,\Delta_{0}\in{\cal R}$,}$ (17)
is the fraction of the number of detections in $\Delta$ over the total number of detections observed in the laboratory $\Delta_{0}$ when an ensemble of identical systems is prepared in the initial state $\rho$}.
We finally consider the general case of $B^{v,V}_{\Delta_{0}}(\Delta)$ with $V\neq I$. Suppose that $V:{{\cal H}}\to{{\cal H}}$ is unitary and, informally speaking,
$B^{v,V}_{\Delta_{0}}(\Delta)=V\frac{1}{\sqrt{A^{v}(\Delta_{0})}}A^{v}(\Delta)\frac{1}{\sqrt{A^{v}(\Delta_{0})}}V^{\dagger}$ (18)
whose rigorous interpretation is (8) if we only know that $Ker(A^{v}(\Delta_{0}))=\{0\}$ without strictly positive lower bounds for $A^{v}(\Delta_{0})$. The spectral functional calculus permits us to rewrite the identity above equivalently as
$B^{v,V}_{\Delta_{0}}(\Delta)=\frac{1}{\sqrt{VA^{v}(\Delta_{0})V^{\dagger}}}VA^{v}(\Delta)V^{\dagger}\frac{1}{\sqrt{VA^{v}(\Delta_{0})V^{\dagger}}}=B^{\prime v}_{\Delta_{0}}(\Delta)$
where, obviously,
$B^{\prime v}_{\Delta_{0}}(\Delta)=\frac{1}{\sqrt{A^{\prime v}(\Delta_{0})}}A^{\prime v}(\Delta)\frac{1}{\sqrt{A^{\prime v}(\Delta_{0})}}\quad\mbox{with}\quad A^{\prime v}(\Delta_{0}):=VA^{v}(\Delta_{0})V^{\dagger}\>.$
Now $({{\cal H}},{\cal R},A^{\prime},U^{\prime})$ is a (causal) relativistic spatial localization observable if $({{\cal H}},{\cal R},A,U)$ is one – where $U^{\prime}_{g}:=V^{\dagger}U_{g}V$. Furthermore the time-translation selfadjoint generators are bounded from below for the former relativistic spatial localization observable if and only the ones of the latter are bounded from below. Hence the apparently more general POVM $B^{v,V}_{\Delta_{0}}(\Delta)$ is still a conditional POVM and simply boils down to a redefinition of the initially used relativistic spatial localization observable. In spite of this straightforward mathematical result we stress that this redefinition may have physical relevance.
5.4 Counterfactual interpretation of $B_{\Delta_{0}}(\Delta)$, and generalization
In establishing the fundamental estimate (17), we assumed that measurements are performed with a relativistic localization observable $A$ whose detectors are also distributed outside $\Delta_{0}$, on the whole rest space $\Sigma\supset\Delta_{0}$, and not only in $\Delta_{0}$. However, it seems physically plausible that $tr(\rho B_{\Delta_{0}}(\Delta))$ coincides with the said conditional probabilities, also in experiments where the detectors of $A$ are switched off outside $\Delta_{0}$, or are not even placed on $\Sigma\setminus\Delta_{0}$. This assumption is similar to other counterfactual interpretations used also in elementary physics [^17] Think of the operational definition of the electric field $\vec{E}$ in terms of its action on a test charge which is not present.. We stress that we are not saying that $A$ does not exist; we are merely saying that the corresponding detectors are placed only in $\Delta_{0}$, so that detection events are registered there only. In support of this counterfactual interpretation of $B_{\Delta_{0}}$, we stress that the fraction on the right-hand side of (17) can be experimentally computed by referring only to the frequencies of events detected in $\Delta_{0}$ (even if, if $\Sigma$ is filled with detectors, there are also events outside $\Delta_{0}$). Furthermore, if two relativistic spatial localization observables $A$ and $A^{\prime}$ coincide on $\Delta_{0}$, they give rise to the same fraction (17) and the same trace $tr(B_{\Delta_{0}}(\Delta)\rho_{\Delta_{0}}^{I})$ for any given state $\rho$. There is another point in support of this interpretation. As we proved in [DRMo24], localization observables can be extended to generic non-flat Cauchy surfaces $S$ (and in [CDM26] this result has been further generalized to maximal achronal sets). However, if $\Delta_{0}$ is common to a pair of Cauchy surfaces $S,S^{\prime}\supset\Delta_{0}$, the corresponding effects coincide: $A_{S}(\Delta_{0})=A_{S^{\prime}}(\Delta_{0})$, and thus so do the corresponding fractions (17). Furthermore, from the definitions of $B_{\Delta_{0}}(\Delta)$, this coherence property makes $tr(B_{\Delta_{0}}(\Delta)\rho)$ independent of $S$ and $S^{\prime}$ when, obviously, also $\Delta\subset\Delta_{0}\subset S\cap S^{\prime}$. In summary, the above counterfactual interpretation can be rephrased by stating that the POVM $B_{\Delta_{0}}$ for $\Delta\subset\Delta_{0}$ measures conditional probability [^18] For states with probability of finding the particle in $\Delta_{0}$ close to $1$. to detect the system in the laboratory based on $\Delta_{0}$. This is true independently of (a) the existence of other detectors outside the laboratory and (b) the choice of other instruments outside $\Delta_{0}$ (in its causal completion) which may give rise to a complete relativistic spatial localization in the whole spacetime and agree with the instruments in $\Delta_{0}$.
[…]
$\Delta_{0}\supset\Delta\mapsto B^{v}_{\Delta_{0}}(\Delta):=\frac{1}{\sqrt{T^{v}(\Delta_{0})}}T^{v}(\Delta)\frac{1}{\sqrt{T^{v}(\Delta_{0})}}\>,$ (19)
[…]
The appearance of unitary operators $V$
$\Delta_{0}\supset\Delta\mapsto B^{v}_{\Delta_{0}}(\Delta):=V\frac{1}{\sqrt{T^{v}(\Delta_{0})}}T^{v}(\Delta)\frac{1}{\sqrt{T^{v}(\Delta_{0})}}V^{\dagger}\>,$ (20)
does not change the structure since (20) can be rewritten as in (19) simply by referring to the non-normalized POVM ${\cal R}\ni\Delta\mapsto T^{\prime v}(\Delta):=VT^{v}(\Delta)V^{\dagger}\in{\mathfrak{B}}({{\cal H}})$.
5.5 Possible commutativity of local conditional effects
We focus on the scalar case only, without loss of generality. Consider a pair of laboratories determined by corresponding non-empty Borel sets $\Delta_{0}\subset\Sigma$ and $\Delta_{0}^{\prime}\subset\Sigma^{\prime}$ (assumed to be open and bounded). We also suppose we have two conditional POVMs as above, $\mathscr{B}(\Delta_{0})\ni\Delta\mapsto B_{\Delta_{0}}(\Delta)\in{\mathfrak{B}}({{\cal H}})$ and $\mathscr{B}(\Delta^{\prime}_{0})\ni\Delta^{\prime}\mapsto B_{\Delta^{\prime}_{0}}(\Delta^{\prime})\in{\mathfrak{B}}({{\cal H}})$. It seems natural to assume that LDP is also valid for the considered POVMs. At this point, the same argument leading to Proposition 4.3, noticing that now the POVMs are normalized in the respective finite sets $\Delta_{0}$ and $\Delta_{0}^{\prime}$, yields, for $\Delta\subset\Delta_{0}$ and assuming that $\{B_{\Delta_{0}}(\Delta),I-B_{\Delta_{0}}(\Delta)\}$ is localized in ${\cal O}_{\Delta}$, namely ${\cal O}_{\Delta}\supset\Delta_{0}$, instead of the much more demanding condition ${\cal O}_{\Delta}\supset\Sigma$ that we found when dealing with the effects $A(\Delta)$. Similarly, if $\Delta^{\prime}\subset\Delta^{\prime}_{0}$ and $\{B_{\Delta_{0}}(\Delta^{\prime}),I-B_{\Delta^{\prime}_{0}}(\Delta^{\prime})\}$ is localized in ${\cal O}_{\Delta^{\prime}}$, then ${\cal O}_{\Delta^{\prime}}\supset\Delta^{\prime}_{0}$. In the absence of obstructions of a different nature which may arise from the specific structure of the considered operators, if $\Delta_{0}$ and $\Delta_{0}^{\prime}$ are sufficiently far apart and causally separated, it should be possible to choose the open sets ${\cal O}_{\Delta}$ and ${\cal O}_{\Delta^{\prime}}$ causally separated. At this point, NSC or RCC would imply that $B_{\Delta_{0}}(\Delta),B_{\Delta^{\prime}_{0}}(\Delta^{\prime})]=0$ through Theorem 4.1. With a more sophisticated discussion relying upon Remark 4.2, the same result could be established by exploiting Beck’s Theorem 3.9.
We stress that this is not a proof of commutativity, since we did not prove that ${\cal O}_{\Delta}$ and ${\cal O}_{\Delta^{\prime}}$ can be chosen causally separated! Different types of issues may arise. However, the problem leading to Corollary 4.4 does not occur here, because the (normalized) POVMs are now defined only on $\Delta_{0}$ and $\Delta^{\prime}_{0}$ and they measure conditional probabilities. For instance, $B_{\Delta_{0}}(\Delta)$ measures the probability of finding the particle in $\Delta\subset\Delta_{0}$, conditioned on the fact that we already know that the particle was found in $\Delta_{0}$ and thus not in $\Delta^{\prime}_{0}$. Hence it is not related to the analogous conditional probabilities of events in $\Delta^{\prime}\subset\Delta^{\prime}_{0}$ measured by $B_{\Delta_{0}^{\prime}}(\Delta^{\prime})$, which considers a disjoint family of events. The price to pay is that this direct physical meaning of these POVMs can be given only when dealing with states $\rho$ for which the probability of detecting the system outside the corresponding laboratory is negligible.
Another technical point in favor of possible commutativity of the effects $B_{\Delta_{0}}(\Delta)$ and $B_{\Delta^{\prime}_{0}}(\Delta^{\prime})$ (for $\Delta_{0}$ and $\Delta^{\prime}_{0}$ causally separated) is the following observation. If we explicitly start from a relativistic spatial localization observable $A$, the family of effects $B_{\Delta_{0}}(\Delta)$, varying $\Delta_{0}$ in the set of bounded Borel sets on all spatial $3$-planes $\Sigma$ such that $Ker(A(\Delta_{0}))=\{0\}$ and $\Delta\in\mathscr{B}(\Delta_{0})$, does not satisfy the hypotheses of Halvorson-Clifton’s Theorem [HaCl02], even if the energy-boundedness requirement is satisfied. This is because the additivity condition 1 fails: additivity is not valid across different laboratories. A more complicated relation holds. If $\Delta_{0}\cap\Delta_{0}^{\prime}=\varnothing$ and $\Delta_{0},\Delta^{\prime}_{0}\subset\Sigma$,
$B_{\Delta_{0}\cup\Delta_{0}^{\prime}}(\Delta)\psi=\frac{1}{\sqrt{A(\Delta_{0}\cup\Delta_{0}^{\prime})}}\left(\sqrt{A(\Delta_{0})}B_{\Delta_{0}}(\Delta\cap\Delta_{0})\sqrt{A(\Delta_{0})}\right$.
$\left.\sqrt{A(\Delta_{0})}B_{\Delta^{\prime}_{0}}(\Delta\cap\Delta^{\prime}_{0})\sqrt{A(\Delta^{\prime}_{0})}\right)\frac{1}{\sqrt{A(\Delta_{0}\cup\Delta_{0}^{\prime})}}\psi$
for $\Delta\subset\Delta_{0}\cup\Delta_{0}^{\prime}$ and for $\psi$ in the natural domain of the composition of operators on the right-hand side.
In a subsequent paper, we will show that conditional POVMs describing spatial localization in laboratories exist and satisfy the commutativity requirement as stated above. The construction will be presented in the spirit of the rigorous Araki-Haag-Kastler formulation of (free) QFT in Minkowski spacetime for a massive real scalar field.
We emphasize that, if $B_{\Delta_{0}}(\Delta)$ belongs to a local algebra in the sense of AHK, then $B_{\Delta_{0}}(\Delta)$ should arise as the restriction to the one-particle space of an operator defined on the full Hilbert space, typically a Fock space of a local quantum field theory on Minkowski spacetime with vacuum state $\Omega$. As already observed, the Reeh–Schlieder theorem then implies that positivity of $B_{\Delta_{0}}(\Delta)$ on the full Hilbert space is incompatible with the requirement that the $\langle\Omega|B_{\Delta_{0}}(\Delta)\Omega\rangle$ be zero, in other words a dark count phenomenon occurs, completely due to the properties of the vacuum state. This is a well-known issue in local quantum physics, recently re-examined in a quantitative framework in [FaCo26] by considering a general model of detector and proving a theoretical inequality to compare the detector efficience with the dark count phenomenon. In that work, the authors also raise the question whether the localization spacetime region of the detector represented by $B_{\Delta_{0}}(\Delta)$ is genuinely concentrated around $\Delta$, or rather should be understood as a larger region encompassing the whole laboratory, as suggested by our perspective.
In the scattering-theory literature based on local QFT, by contrast, a different standpoint is usually adopted: detector operators are assumed to assign the value zero to the vacuum state: $\langle\Omega|B\Omega\rangle=0$, and operators $B$ are therefore not local (but quasi-local as is the case for the Haag–Ruelle scattering theory [Haa96, Ara09]) contrarily to the operators we described above.
6 Conclusions and outlook
In this work, we studied the issue of commutativity in the representation of relativistic locality within the description of localization observables for relativistic quantum systems. The celebrated no-go theorem of Halvorson and Clifton shows that the commutativity of causally separated localization effects is incompatible with other apparently natural assumptions concerning the notion of spatial localization. This implies, in particular, that localization operators – elements of suitable families of POVMs associated with the rest spaces of inertial observers – cannot belong to local algebras of observables in the sense of Araki–Haag–Kastler. Adopting Busch’s analysis of commutativity, this paper established, in particular, that the commutativity requirement is not, in this context, a necessary consequence of more elementary principles such as no-signaling or relativistic consistency. This holds for elementary systems, i.e., particles, as soon as one assumes that, in highly idealized experiments in which their position is searched for throughout the whole rest space of a reference frame at a given time, they have the propensity to localize at a unique place. In this sense, in the case under consideration, the failure of commutativity does not represent a problem for causality.
On the other hand, we proved that, in principle, both commutativity and localization can be restored when one refers to a more sophisticated, and perhaps more realistic, notion of conditional localization. Here one tries to define POVMs normalized on a restricted region of space, namely a laboratory. These POVMs correspond to the conditional probability of finding a particle in a subregion of the laboratory, given that it has been detected there. A formal expression for this type of POVM, for states of a particle with small probability of being detected outside the laboratory, has been suggested in terms of general localization observables associated with full rest spaces. It is argued that effects belonging to pairs of these conditional POVMs, associated with causally separated laboratories, may in principle commute. A crucial technical result in this interpretation is the so-called gentle measurement lemma.
It would be interesting to explore possible connections between these ideas and the modular-theoretic approach developed in [LedO26]. This issue will be investigated elsewhere.
As a final remark, we observe that, in light of the results presented here – and also considering the failure to describe localization by means of a self-adjoint operator – we may conclude that the notion of position for a physical object in a fully developed relativistic quantum theory has a rather peculiar status. In contrast with non-relativistic quantum theory, there does not appear to be a preferred notion of position observables. Although one expects that, in the non-relativistic regime, all such notions converge to the naive one – namely, the position operator (as shown, in particular, for the relativistic localization observables constructed in [Mor23, DRMo24] – this is no longer the case in the fully relativistic setting. Furthermore, any notion of spatial localization must necessarily be unsharp. Position in a given reference frame cannot be identified with the joint spectrum of a triple of self-adjoint operators. From this perspective, unsharp observables assume a foundational role and should not be regarded merely as approximate or effective versions of sharp ones. It is plausible that sharp observables, i.e., self-adjoint operators, are instead to be interpreted primarily as generators of symmetries.
In a forthcoming paper, we shall construct explicit examples of general localization observables and conditional local position observables using the local AHK approach to QFT. In particular, we shall see that this type of localization observable can be constructed from the (smeared) normally ordered stress–energy operator, taking advantage of some known results on quantum energy inequalities.
Acknowledgments
The author is grateful to D. Castrigiano and C. De Rosa for useful discussions about the issues presented in this work. This work has been written within the activities of INdAM-GNFM.
Appendix
Appendix A Appendix: H-C’s hypothesis (4) validity for some fermionic PVMs
According to CC, $A(\Delta)\leq A(J^{+}(\Delta)\cap(\Sigma+t{a}))\quad\mbox{if $t>0$.}$ Since the effects are actually orthogonal projectors, that inequality is equivalent to (e.g., see [Mor19])
$A(\Delta)A(J^{+}(\Delta)\cap(\Sigma+t{a}))=A(J^{+}(\Delta)\cap(\Sigma+t{a}))A(\Delta)=A(\Delta).$
On the other hand, if the distance in $\Sigma$ between $\Delta$ and $\Delta^{\prime}$ is strictly positive, then
$(\Delta^{\prime}+t{a})\cap J^{+}(\Delta)=\varnothing\quad\mbox{if $t>0$ is sufficiently small,}$
so that $A(J^{+}(\Delta)\cap(\Sigma+t{}))A(\Delta^{\prime}+t{a})=0$. Finally,
$A(\Delta)A(\Delta^{\prime}+t{a})=(A(J^{+}(\Delta)\cap(\Sigma+t{a}))A(\Delta))A(\Delta^{\prime}+t{a})$
$A(\Delta)A(J^{+}(\Delta)\cap(\Sigma+t{a}))A(\Delta^{\prime}+t{a})=0\>.$
In particular, $A(\Delta),A(\Delta^{\prime}+t{a})]=A(\Delta)A(\Delta^{\prime}+t{a})-A(\Delta^{\prime}+t{a})A(\Delta)$
$A(\Delta)A(\Delta^{\prime}+t{a})-(A(\Delta)A(\Delta^{\prime}+t{a}))^{\dagger}=0.$
The case $t<0$ is analogous.